### v3

# Large Language Model-Driven Closed-Loop UAV Operation with Semantic Observations

Abstract
Recent advances in large Language Models (LLMs) have revolutionized mobile robots, including unmanned aerial vehicles (UAVs), enabling their intelligent operation within Internet of Things (IoT) ecosystems. However, LLMs still face challenges from logical reasoning and complex decision-making, leading to concerns about the reliability of LLM-driven UAV operations in IoT applications. In this paper, we propose a LLM-driven closed-loop control framework that enables reliable UAV operations powered by effective feedback and refinement using two LLM modules, i.e., a Code Generator and an Evaluator. Our framework transforms numerical state observations from UAV operations into natural language trajectory descriptions to enhance the evaluator LLM’s understanding of UAV dynamics for precise feedback generation. Our framework also enables a simulation-based refinement process, and hence eliminates the risks to physical UAVs caused by incorrect code execution during the refinement. Extensive experiments on UAV control tasks with different complexities are conducted. The experimental results show that our framework can achieve reliable UAV operations using LLMs, which significantly outperforms baseline approaches in terms of success rate and completeness with the increase of task complexity.
Index Terms:
UAV, Large Language Model, Robotics, IoT.
I Introduction
Uavs, or drones, are increasingly being integrated into IoT ecosystems to support various applications, such as surveillance and monitoring, smart cities, and infrastructure inspection [1]. Their high mobility and rich sensing capabilities make them an ideal candidate for data collection, communication, and service delivery tasks in these applications. Designing UAVs that have intelligence to perform complex tasks in these applications correctly remains challenging, as it demands both robust instruction interpretation and executing tasks with robust reasoning. Recent advances in LLMs [2, 3, 4, 5] have demonstrated remarkable proficiency in robotic areas such as planning [6], navigation [7, 8], communication [9], and automation [10]. With LLMs’ exceptional natural language (NL) understanding and context generation capabilities, LLMs offer a promising avenue for robots with greater intelligence. Motivated by these developments, research efforts have been made towards the reliable integration of LLMs and UAVs [11, 12, 13, 14, 15], to enable LLM-driven UAV operations.
Despite LLMs’ potential for various robotic tasks, LLMs continue to face challenges from logical reasoning [16], undermining their reliability in tasks that require complex decision-making. Crucially, unlike text generation applications where semantic-level accuracy is sufficient, UAV operation requires the precise execution of sequential actions. Erroneous or invalid control code can lead to unsafe UAV behavior and pose safety and security risks to the public, such as UAV crashes. Recent studies leverage prompt engineering [17] to enhance LLM in translating NL instructions into control code [11, 18, 15]. However, these approaches adopt open-loop designs and lack mechanisms to verify the correctness of the synthesized code, leaving questions such as cumulative error and temporal action consistency unsolved [19].
Figure 1: Illustration of LLM-driven closed-loop UAV Operation and Refinement.
Inspired by control theory, more recent studies have adopted a closed-loop architecture that enables evaluation of generated code and provides feedback for refinement, thereby steering robot execution towards prescribed task objectives [20, 21, 22, 23, 24, 25, 26]. Nevertheless, challenges are prone. First, existing approaches prompt LLM with numerical state observations in scenarios that require insights into robot dynamics for evaluation [24, 26]. Given that LLMs exhibit limited numerical reasoning and processing capabilities [27], an LLM-driven evaluation may overlook the latent semantics of robot actions and fail to achieve a comprehensive assessment of the robot’s movements and dynamics, leading to unreliable code refinement feedback. Second, current closed-loop approaches require physical robot execution to generate state observations for evaluation [22, 24]. Such an approach may be acceptable for ground robots where occasional errors rarely cause irreparable damage. However, it poses risks for aerial platforms, as even slight control deviations can precipitate UAV crashes.
To address these challenges, we propose a novel closed-loop feedback and refinement framework that leverages semantic observations to enhance the reliability of LLM-driven UAV control as depicted in Fig. 1. Our framework comprises two LLM modules: a Code Generator, which synthesizes or refines UAV operation code based on task description or evaluation feedback; and an Evaluator, which assesses code performance and provides feedback to guide code refinement. The process unfolds as follows: 1) The code generator first produces an initial code based on the task description; 2) The code is then execute in simulation to produce raw numerical state observation, which are subsequently transformed into an NL trajectory observation; 3) The evaluator compares NL trajectory observation against the original task specification and returns evaluation feedback; 4) If the evaluation confirms the trajectory meets the task objectives, the code will deploy on the UAV for mission execution. Otherwise, the evaluation feedback is returned to the code generator for iterative refinement. By introducing NL trajectory representations rather than numerical state information, our method significantly enhances the evaluator’s comprehension of the UAV’s state dynamics, thereby improving the reliability and precision of the evaluation feedback. Moreover, our method enables the refinement process in simulation, which eliminates the risks to UAVs caused by incorrect code execution during refinement, and hence supports safe and efficient development.
We conduct extensive experiments on UAV operation tasks with varying levels of complexity to evaluate the effectiveness of our proposed framework. The experimental results show that our framework achieves a $100\%$ success rate and $100\%$ completeness for basic operations. For complex tasks, our framework achieves an $85\%$ success rate and $98.5\%$ completeness. Compared with baseline approaches [15, 25, 24], our framework greatly improves both success rate and completeness when the task complexity increases. The contributions of our work can be summarized as follows:
  • Closed-Loop Refinement: A novel closed-loop feedback and refinement framework with a code generator and an evaluator designed to enhance the reliability of LLM-driven UAV control.
  • Semantic Encoding: A novel algorithm for converting numerical state observations into step-wise NL trajectory descriptions, thereby enhancing the evaluative reasoning process.
  • Safe Development: A simulation-based LLM-generated UAV code evaluation that replaces physical deployment to prevent potential UAV crashes in code refinement.
  • Comprehensive Evaluation: A comprehensive experimental evaluation that demonstrates the performance of our framework with a detailed analysis of design factors.
The remainder of this paper is structured as follows: Section II reviews and discusses existing research. Section III presents our problem formulation, which is followed by the detailed design of our proposed framework in Section IV. Section IV-D provides the evaluation and analysis of our framework. This paper is concluded in Section VI.
II Related Work
II-A LLM for Robotics
LLMs have demonstrated exceptional capabilities in semantic understanding and context generation, making them invaluable in robotics. ChatGPT for Robotics [11] and Code as Policies [18] have shown significant progress in LLM to control various types of robots. In the domain of aerial platforms, TypeFly [12] introduces an end-to-end LLM framework for piloting drones via human languages, while REAL [28] integrates LLMs into a mission planning and control framework. Moreover, continual advancements in LLMs’ reasoning and code generation capabilities [5, 29] are opening new opportunities for a variety of LLM-driven robotic tasks. Despite these advances, robotic performance continues to be limited by the reasoning capacity of LLMs, which frequently demonstrate unreliability when faced with tasks requiring complex reasoning [11].
II-B Prompt Engineering
Recent studies have increasingly employed prompt engineering [17] to improve the reliability of LLM-driven robotic systems. For instance, GSCE proposes a prompt framework as a system prompt to improve the reliability of LLM-driven UAV control code generation [15]. Prompt engineering typically leverages two core techniques: in-context learning [30] and Chain-of-Thought (CoT) reasoning [31]. In-context learning enables the LLM to identify patterns from a limited set of examples and apply them to new situations. This facilitates the generation of control code that adheres to robot policy and how to ground task descriptions from a few examples [18, 32]. On the other hand, CoT prompts the LLM to articulate intermediate inference steps, which is valuable for robotic tasks requiring sequential, stepwise decision-making. By decomposing a task into steps, CoT encourages the production of code that aligns with each stage of the intended action plan. Prior studies have embedded CoT within the examples to guide LLMs through reasoning step-by-step via in-context learning [21, 32]. In this study, we integrate both CoT and in-context learning strategies to enhance the quality of code generation and evaluation in LLM-driven UAV control.
II-C LLM-based Feedback, Evaluation, and Refinement
ChatGPT for Robotics emphasizes the necessity of human oversight for code generation [11], reflecting the fact that robotic performance remains constrained by LLMs’ reasoning limitations. Moreover, LLMs exhibit sensitivity to input perturbations [33, 34, 35]. Thus, recent research efforts adopt closed-loop refinement strategies to verify the LLM generation for improving the reliability of LLM-driven control. For instance, PluginSafety [36] designs a module that enforces predefined logical constraints evaluation to ensure LLM-generated content adheres to predefined constraints. More recent studies employ corrective feedback to iteratively refine generated code. Some integrate human or automated checker feedback [20, 21, 37, 38] to re-prompt the LLM and improve reasoning accuracy. While others monitor the robot status during execution and provide feedback, such as runtime errors, robot constraints violations [39, 22, 23, 24]. The evaluation processes generally fall into two categories: self-evaluation and external evaluation. SELF-REFINE [25, 39] leverages a single LLM to assess and refine its own outputs based on feedback, whereas external evaluation approaches employ a secondary LLM or external evaluator to generate corrective feedback [36, 22, 23, 24]. In our approach, numerical state information is first transformed into NL observations, and then we use an external LLM-evaluator to provide feedback for iterative refinement.
Figure 2: An example of how our LLM-driven closed-loop feedback and refinement framework works during solving a task that requires examining a square area. In the first loop, the evaluator identifies deviations in the NL trajectory observation (action 12), providing error feedback into the code generator to refine the code. In the second loop, the evaluator confirms refined code matches the task, and the code is deployed on the UAV for mission execution.
III Problem Formulation
Given a UAV operation task description, the framework proposed in this paper aims to use LLMs to reason through the task description and generate accurate operation code after observation, evaluation, feedback, and refinement. Specifically, we formulate our problem by extending the LLM-driven UAV operation problem in [15] as $P=\langle s_{0},S,A,T,G\rangle$, where $s_{0}$ is the initial state, $S=(s_{0},s_{1},s_{2},\dots,s_{n})$ is the finite set of discrete states encoding the UAV’s spatial position and orientation, $A$ is the finite action (UAV skills) set, $T$ is the deterministic transition, and $G\subseteq S$ is the goal state. For a current state $S$, an action $a\in A$ is selected from the skill set $A$, leading to a deterministic state transition: $T:S\times A\rightarrow S$ that determines the next state. A candidate solution to problem $P$ is a sequence of actions $l=(a_{1},a_{2},\dots,a_{n})$ that control the UAV from the initial state $s_{0}$ through intermediate states until reaching a final state $s_{n}$. The state transitions of $l$ yield state observations $O$, and evaluation $E=\langle O,P\rangle$ assesses whether $O$ satisfies the problem objectives and outputs $F$ as feedback. If $F$ is ”YES”, the sequence of actions $l$ is deemed a valid solution solution that flies the UAV from $s_{0}$ to the goal state $(s_{n}\equiv G$); otherwise, $l$ is iteratively refined based on the feedback $F$.
IV Methodology
IV-A Overview
The overall idea of our proposed reliable LLM-driven closed-loop control framework for UAV operations is illustrated in Fig. 2 and Algorithm 1. Given a UAV operation task description by the user, the LLM configured as the code generator with enhanced reasoning capabilities towards UAV operation first produces the initial version of the operation code. A simulation-based execution of the code is then performed to obtain the numerical state observations. To close the gap between numerical state observations and the LLM-driven evaluation, a state transformation is designed that transforms numerical state observations into an NL trajectory observation. The evaluator analyzes the trajectory observation together with the task description. If the trajectory satisfies all task objectives in evaluation, the operation code is ready for deployment; otherwise, the evaluator provides feedback to the code generator for refinement. This process iterates until the task objectives are achieved or the maximum number of iterations is reached. We now provide the detailed design of our code generation, trajectory observation, evaluation, feedback, and refinement in the following sections.
Input: $Task$: Task description
Output: $Code$: Code script
$Code\leftarrow$ Code Generator$(Task)$
$itr\leftarrow 1$
while $itr\leq$ maximum round do
       $O\leftarrow$ State Transformation$(Code)$
      $F\leftarrow$ Evaluator$(O,Task)$
      if “YES” in $F$ then
             break
      else
             $Code\leftarrow$ Code Generator$(F)$
       end if
      $itr\leftarrow$ $itr+1$
end while
return $Code$
Algorithm 1 Closed-loop Feedback and Refinement
IV-B Code Generation
The LLM-driven code generation in our framework is achieved by configuring a general-purpose LLM agent (code generator) with UAV operation-related system prompts. We adopt prompt engineering strategies proposed by the GSCE framework [15], which have been demonstrated to be effective for enhancing the reasoning capabilities of LLMs in basic UAV operation code generation. Specifically, the system prompts comprise guidelines, skill APIs, constraints, and examples. The guidelines shape the LLM’s role for code generation, the skill APIs empower LLM to employ appropriate functions, the constraints regulate LLM with robotic policies, and the examples are pairs of query-response that serve as illustrations. These strategies serve as the basic building blocks to support the LLM-driven UAV operation code generation. On top of that, our code generation also utilizes constructive feedback from our evaluation design in section IV-D. The feedback helps the code generator obtain a more comprehensive understanding of the dynamics and interconnections among UAV operations implied by the task description, which enhances the LLM’s reasoning for code generation and code refinement.
IV-C Trajectory Observation
Trajectory observation constitutes the foundation of our closed-loop feedback. It embeds the observations of the UAV’s historical trajectory dynamics, thereby enabling the evaluator to identify its deviations from the task objectives. Rather than getting trajectory observations from real robots [24, 22] that can cause a UAV crash if the LLM-generated code contains errors, our framework adopts simulation-based execution to generate trajectory observations. When the LLM-generated code is executed in simulation, it produces the UAV’s state observations for each action. These observations are numerical representations of the UAV’s states, including the UAV’s position and orientation. However, numerical observations offer limited semantic insight for an LLM-based evaluator because LLM lacks prior knowledge about interpreting the semantic meaning of numerical state tuples [27].
To bridge this gap, we transform each numerical state observation into a step-wise NL description and aggregate these into an NL trajectory observation. As shown in Fig. 1, NL description captures the semantic meaning of the state changes, such as “Move 5 meters West while facing North, the UAV moves to [0, 5]”. This transformation provides the evaluator with a clear understanding of the semantic significance of each action and overall trajectory dynamics, thereby improving the evaluator’s ability to identify deviations and generate precise feedback for code refinement.
Given the LLM-generated UAV operation code executed in simulation, Algorithm 2 presents the transformation from numerical state observations to the NL trajectory observation $O$. For each action identified in the $Code$, its execution in the simulation will be observed to obtain the corresponding numerical UAV state observations. Specifically, for each action $a$, the simulator first executes $a$, if a runtime error occurs, the algorithm collect and log errors alongside the corresponding action identifier in $O$; otherwise, the algorithm computes the difference between the current state $s_{\mathrm{current}}$ and the last state $s_{\mathrm{last}}$, denoted as $\Delta s=s_{\mathrm{current}}-s_{\mathrm{last}}$. For each element in the state difference $\Delta s$, the algorithm appends the corresponding NL description of orientation and changes in yaw, x-axis, y-axis, or z-axis to $O$. The algorithm also appends the UAV’s NED position to $O$. Upon completion of all actions, the algorithm returns the complete NL trajectory observation $O$.
Input: $Code$: Code script
Output: $O$: Trajectory observation
$O\leftarrow\emptyset$
foreach  $action$ in $Code$  do
       $s_{\mathrm{last}}\leftarrow$ State Observation $(x,y,z,\theta)$
      Execute $action$
      if error executing $action$ then
             $O\leftarrow O$ + “Error in executing $action$ with error message $error$
       end if
      $s_{\mathrm{current}}\leftarrow$ State Observation $(x,y,z,\theta)$
      $\Delta s\leftarrow s_{\mathrm{current}}-s_{\mathrm{last}}$
      foreach $element$ in $\Delta s$ do
             if $element$ in orientation then
                   $O\leftarrow O$ + Rotate $element$ degrees in Yaw.
            else
                  if $element$ in NED axis then
                         $O\leftarrow O$ + “Move $element$ meters in North, East, or Down while facing $s_{\mathrm{current}}[\theta]$.”
                   end if
             end if
            $O\leftarrow O$ + “The UAV moves to $s_{\mathrm{current}}[x,y,z]$.”
       end foreach
end foreach
return $O$
Algorithm 2 State Transformation
IV-D Evaluation
In our framework, we design the evaluation by leveraging the LLM’s semantic reasoning capabilities to analyze the NL trajectory observation. The evaluator aims to identify deviations (if any) between the NL trajectory observation and the task description, and then construct feedback based on the deviations. Instead of using the same LLM used for code generation for self-feedback evaluation, our design employs a separate LLM to prevent the same LLM from amplifying its bias [40].
To improve the precision of LLM’s evaluation, we adopt the design principles of the code generator by configuring the evaluator as an LLM agent with a structured system prompt. However, unlike the code generator, the evaluator does not produce code but instead generates evaluation feedback. To address this issue, we construct the evaluator’s system prompt with roles, rules, and references, as depicted in Fig. 3. Specifically, the roles define the LLM agent as an evaluator that compares the NL trajectory observation against the task description, while the rules serve as explicit evaluation criteria to direct the LLM when reasoning about trajectories. Additionally, the references are exemplar trajectories that serve as benchmarks to support the evaluation of NL trajectory observations. Furthermore, the system prompt requires the evaluation outcome to be informative and constructive by providing explanations about the depicted deviations. As demonstrated in Fig. 2, given a NL trajectory observation and a task description, the structured system prompt enables the LLM-evaluator to identify deviations from the task description and explain the errors in the trajectory observation (e.g., “Action 12 flies West instead of East”). Consequently, the code generator is aware of the errors and could correctly refine the code in the subsequent refinement process.
Figure 3: The evaluation process begins by configuring the LLM agent (evaluator) via a system prompt framework that includes roles, rules, and references. Next, the trajectory observations and the task description are provided to the evaluator for comparison.
IV-E Feedback and Refinement
The feedback and refinement steer the LLM-generated UAV operation code towards fully satisfying the task objectives. They leverage the two LLM agents designed in section IV-B and section IV-D and form a closed-loop as illustrated in Fig. 2. Specifically, the code generator interprets the evaluation feedback $F$ and refines the generated code accordingly. Then, the refined code is executed in the simulation again to produce a new NL trajectory observation $O^{\prime}$ for re-evaluation. The feedback and refinement loop continues iteratively until the evaluator outputs ”YES”, indicating that $E$ confirms that the refined action sequence $l$ successfully solves the task problem $P$ and the UAV operation code is ready for deployment. In addition, if the maximum number of iterations is reached without a successful evaluation, the evaluator suggests human intervention.
V Experiment
The goal of this experiment is to validate the effectiveness of our proposed LLM-driven closed-loop framework in enhancing the reliability of LLM-driven UAV operation code generation. To achieve this, the experiment is designed with three primary goals: (1) to evaluate the performance of our closed-loop feedback and refinement framework in LLM-driven UAV tasks; (2) to analyze the impact of different design choices on the performance of our framework; (3) to assess the behavior of our framework under diverse task scenarios.
V-A Experimental Setup
V-A1 Environment
To ensure safe code execution without risking real-world failures, we conduct experiments in the “AirSim” simulator [41], which realistically simulates UAV (e.g., a quadcopter) dynamics and environmental interactions. During the experiment, UAV state information was fully accessible from the simulator in real time and utilized for closed-loop feedback and evaluation purposes.
V-A2 Tasks
We propose Basic and Advanced task sets to rigorously evaluate the performance of LLM-driven control code generation across varying levels of operational complexity [^1] The complete list of tasks is available at https://github.com/ai-uavsec/CLGSCE. The Basic set comprises 44 tasks adopted from GSCE [15]. Each task involves one or two intuitive discrete movements to evaluate LLM’s few-stage planning and basic spatial reasoning (e.g., “Turn 90° clockwise, then fly 4 meters left in the drone’s body frame”). In contrast, the Advanced set consists of 20 tasks that emulate real-world UAV deployment (e.g., “Take off and fly up 5 meters. You should fly in a square pattern with 5-meter sides by moving forward, right, backward, and left. Make sure the drone is oriented to the flying direction.”). These tasks demand reasoning about the objectives in the task description and require 6-19 UAV actions to reach the goal state. The advanced task set is intended to evaluate LLM’s multi-stage planning and complex spatial reasoning. Additionally, all tasks in both task sets were manually validated to avoid potential simulator-induced errors that could affect the evaluation process.
V-A3 Compared Methods
Three recently proposed methods [15, 25, 24] are selected as the baselines for comparison with our approach, including
  • GSCE: An open-loop approach in which the GSCE code generator synthesizes control code based on the task prompt, without any refinement. [15]
  • LLM Self-feedback (Self-Refine): A single LLM performs both code generation and self-refinement. The LLM analyzes NL trajectory observations to detect deviations from the task and refines code based on its self-assessment.  [25]
  • Feedback with Numerical State Observations (Numerical): This method introduces a secondary LLM serving as evaluator designed according to section IV-D. The evaluator detects deviations from numerical state observations and provides corrective guidance for the code generator for refinement. [24]
All methods described above employ “o3-mini” (o3-mini-2025-01-31) [29] as the foundational LLM. The generated code was then deployed on a quadcopter in AirSim [41] to evaluate the performance of each method. To mitigate the randomness of LLM’s output [42], we repeat each task three times in the experiment.
V-B Evaluation Setup
V-B1 Metrics
The performance is evaluated using two metrics as GSCE [15]: Completeness and Success Rate (SR).
Completeness quantifies the proportion of goal states successfully reached, calculated as the ratio of the total count of correct actions to the total number of required actions of the ground truth trajectory in each task. The completeness offering insights into performance across the intermediate control process.
$Completeness=\frac{\sum\text{correct actions}}{\mathrm{len}(l)}$ (1)
SR is grounded in real-world operational requirements, wherein the UAV must fly the trajectory precisely as humans instructed. As formalized in Eq. 2, a task is counted as a success only when the UAV reaches the final state while following the correct sequence of actions that lead to the ground truth trajectory.
$Success=\begin{cases}1,&Completeness=1,\\ 0,&Completeness<1.\end{cases}$ (2)
V-B2 Ground Truth
The ground truth is represented by a list of state transitions. It is a vector of four elements: $[x,y,z,\theta]$, where $x$, $y$, and $z$ denote the UAV’s position changes in the North, East, and Down axes, and $\theta$ represents yaw rotation. We assume the UAV’s state is fully observable at all times, thereby ensuring that each transition can be precisely monitored and compared against ground truth.
V-C Result and Analysis
V-C1 Overall Result
The average SR and completeness across Basic and Advanced tasks are summarized in Table I. For the Basic tasks, all methods showed remarkable performance. Specifically, the GSCE, Self-Refine, and Ours methods achieved $100\%$ in both SR and completeness, while the Numerical achieved $95.5\%$ SR and $97.3\%$ completeness. This demonstrates their ability to generate correct code for a few-stage planning and basic spatial reasoning. However, the slightly lower performance of the Numerical method suggests that relying on numerical state observations can occasionally lead to misinterpretation in evaluation, resulting in incomplete or incorrect refinements.
Performance on the Advanced tasks declines notably across all methods as a result of increased task complexity. The open-loop GSCE method achieves only $66.7\%$ SR, indicating that refinement becomes critical as task complexity increases. The Self-Refine method performs worse, with $50.0\%$ SR and $74.3\%$ completeness, proving that relying on a single LLM for both generation and self-evaluation amplifies bias [40], resulting in introducing additional errors in refinement. The Numerical method achieves $73.3\%$ SR and $92.4\%$ completeness, showing that it can detect some deviations; however, its numeric perspective still fails to provide reliable evaluation feedback for refinement. In contrast, our proposed method substantially outperforms all baselines, achieving $85.0\%$ SR and $98.5\%$ completeness on Advanced tasks. The overall results demonstrate that our proposed framework significantly enhances the reliability of LLM-driven UAV operation code generation, particularly for complex tasks that require multi-stage planning and complex spatial reasoning.
TABLE I: Average SR and Completeness of Proposed Methods
Base Task Advanced Task SR Completeness SR Completeness GSCE [15] $100.0\%$ $100.0\%$ $66.7\%$ $88.9\%$ Self-Refine [25] $100.0\%$ $100.0\%$ $50.0\%$ $74.3\%$ Numerical [24] $95.5\%$ $97.3\%$ $73.3\%$ $92.4\%$ Ours $100.0\%$ $100.0\%$ $\mathbf{85.0\%}$ $\mathbf{98.5\%}$
V-C2 Failure Cases Analysis
We further analyze the failure cases (cases that $Success=0$) in the Advanced task set. The analysis focuses on three aspects: (1) the number of failures and average completeness for each failed task ID with the respective ground truth action length (Fig. 4); (2) the average completeness of each failed task ID (Fig. 5); (3) the amount of task that experienced failure and the average completeness of all failed tasks (Table II).
The results in Fig. 4 reveal a clear trend that all methods exhibit a notable increase in failures as task complexity increases (i.e., more likely to fail on tasks with longer action length). Specifically, the baseline GSCE method failed in $11$ out of $20$ tasks ($55\%$ of the Advanced set), with an average completeness of $77.4\%$. This indicates that the GSCE method can reason through a substantial portion of the task correctly, however, it still struggles to maintain full alignment with the objectives of the task. This result highlights the need for improvement in the completeness of failed tasks to reduce overall failure rates. The Self-Refine method exhibited the weakest performance, with failures on $17$ tasks ($85\%$ of Advanced tasks). Moreover, the failures in Self-Refine were the least complete among all methods, demonstrating that performing evaluation and refinement on a single LLM not only fails to correct errors but may also introduce more errors. The Numerical method failed on $9$ tasks ($45\%$ of Advanced tasks) and achieved an average completeness of $69.5\%$ only. These results indicate that numerical state observations are insufficient for reliable code refinement and may even lead to erroneous refinement.
Figure 4: Failure analysis across the four methods on the Advanced task set. The x-axis represents the task IDs that are sorted in ascending order based on the length of the task’s action steps. And the y-axis presents two metrics for each task: (1) the count of failures observed over three repetitions (left y-axis); (2) the length of the ground truth action steps (right y-axis).
Figure 5: Average completeness for each failed task on the Advanced task set over four methods. The smaller the value, the less complete the failed task is.
In contrast, our method achieves the best performance, failures occurred only on $4$ tasks ($20\%$ of Advanced tasks), and the completeness of failed tasks arose at $91.6\%$. Moreover, the completeness in Fig. 5 demonstrates that even when failures occur, our method produces code that remains closely aligned with task objectives. These findings underscore the effectiveness of our method in refining errors and improving the completeness of LLM-driven UAV operation, which contributes to an increased success rate in handling complex UAV operation tasks.
TABLE II: Amount of Task ID Experienced Failure and Average Completeness for All Failed Tasks
Task ID Amount $\downarrow$ Failed Tasks Completeness $\uparrow$
GSCE [15] $11$ $77.4\%$
Self-Refine [25] $17$ $48.5\%$
Numerical [24] $9$ $69.5\%$
Ours $\mathbf{4}$ $\mathbf{91.6\%}$
V-C3 Number of Refinement Iterations
[…]
Figure 6: Curves of average SR (left) and completeness (right) over the number of refinement iterations for the NL method on Advanced tasks.
V-D Evaluator Configuration Analysis
As the precision of the evaluation feedback directly influences the quality of code refinement, we also analyzed how different evaluator configurations affect evaluation precision. The evaluation precision is defined as the ratio of correct evaluations (i.e., correctly identifying whether the trajectory aligns with the task and providing a detailed explanation if deviations are identified) to the total number of evaluations conducted. To perform this analysis, we built a dataset from each Advanced task with one correct NL trajectory observation (Correct O) and one incorrect NL trajectory observation (Incorrect O). As presented in Table III, we tested four configurations of the evaluator system prompt: (1) roles; (2) roles $+$ rules; (3) roles $+$ references; (4) roles $+$ rules $+$ references. Moreover, each configuration was evaluated three times to reduce the randomness of LLM’s output [42].
TABLE III: Evaluator Configurations
Configuration (1) (2) (3) (4)
Roles
Rules

### v4

@@
 However, LLMs still face challenges from logical reasoning and complex decision-making, leading to concerns about the reliability of LLM-driven UAV operations in IoT applications.
-In this paper, we propose a LLM-driven closed-loop control framework that enables reliable UAV operations powered by effective feedback and refinement using two LLM modules, i.e., a Code Generator and an Evaluator.
-Our framework transforms numerical state observations from UAV operations into natural language trajectory descriptions to enhance the evaluator LLM’s understanding of UAV dynamics for precise feedback generation.
+In this paper, we propose a closed-loop LLM-driven UAV operation code generation framework that enables reliable UAV operations powered by effective feedback and refinement using two LLM modules, i.e., a Code Generator and an Evaluator.
+Our framework transforms numerical state observations from UAV operations into semantic trajectory descriptions to enhance the evaluator LLM’s understanding of UAV dynamics for precise feedback generation.
 Our framework also enables a simulation-based refinement process, and hence eliminates the risks to physical UAVs caused by incorrect code execution during the refinement.
 Extensive experiments on UAV control tasks with different complexities are conducted.
-The experimental results show that our framework can achieve reliable UAV operations using LLMs, which significantly outperforms baseline approaches in terms of success rate and completeness with the increase of task complexity.
+The experimental results show that our framework can achieve reliable UAV operations using LLMs, which significantly outperforms baseline methods in terms of success rate and completeness with the increase of task complexity.
 Index Terms:
@@
 I Introduction
-Uavs, or drones, are increasingly being integrated into IoT ecosystems to support various applications, such as surveillance and monitoring, smart cities, and infrastructure inspection [1].
+Uavs, or drones, are increasingly being integrated into IoT ecosystems for various applications, such as surveillance and monitoring, smart cities, and infrastructure inspection [1].
 Their high mobility and rich sensing capabilities make them an ideal candidate for data collection, communication, and service delivery tasks in these applications.
-Designing UAVs that have intelligence to perform complex tasks in these applications correctly remains challenging, as it demands both robust instruction interpretation and executing tasks with robust reasoning.
-Recent advances in LLMs [2, 3, 4, 5] have demonstrated remarkable proficiency in robotic areas such as planning [6], navigation [7, 8], communication [9], and automation [10].
-With LLMs’ exceptional natural language (NL) understanding and context generation capabilities, LLMs offer a promising avenue for robots with greater intelligence.
-Motivated by these developments, research efforts have been made towards the reliable integration of LLMs and UAVs [11, 12, 13, 14, 15], to enable LLM-driven UAV operations.
-Despite LLMs’ potential for various robotic tasks, LLMs continue to face challenges from logical reasoning [16], undermining their reliability in tasks that require complex decision-making.
-Crucially, unlike text generation applications where semantic-level accuracy is sufficient, UAV operation requires the precise execution of sequential actions.
-Erroneous or invalid control code can lead to unsafe UAV behavior and pose safety and security risks to the public, such as UAV crashes.
-Recent studies leverage prompt engineering [17] to enhance LLM in translating NL instructions into control code [11, 18, 15].
-However, these approaches adopt open-loop designs and lack mechanisms to verify the correctness of the synthesized code, leaving questions such as cumulative error and temporal action consistency unsolved [19].
+Given a UAV task, the traditional pipeline requires a specialized engineer to write the operation code based on the task description and then deploy the code on the device for execution, as shown in Fig.
+1(a).
+Such a process requires substantial domain knowledge and can also be time-consuming.
+With LLMs’ exceptional semantic understanding and context generation capabilities, recent research has demonstrated the potential of leveraging LLM to facilitate the operation code generation process for a given robotic task [2, 3, 4, 5, 6, 7, 8, 9, 10, 11].
+As illustrated in Fig.
+1(b), these LLM-driven approaches enable the generation of UAV operation code with natural language tasks, which significantly simplifies the code generation and also brings the potential of automated UAV task deployment.
+Figure 1: Pipelines of robotic operation code generation: (a) without LLMs and (b) with LLMs.
+Despite the benefits brought by LLM-driven UAV operation code generation, it also raises concerns about the reliability of LLM-generated code due to the challenges faced by LLM in terms of logical reasoning and complex decision-making [12].
+Different from text generation applications, where semantic-level accuracy is sufficient, the generation of UAV operation code requires the precise execution of sequential actions.
+Erroneous or invalid operation code can lead to unsafe UAV behavior and pose safety and security risks to the public, such as UAV crashes.
+To enhance the reliability of LLM in translating natural language tasks into operation code, techniques such as prompt engineering [13] and few-shot learning [14] have been widely adopted when designing LLM-driven operation code generation research for robots, including UAVs, and have demonstrated promising results.
+However, a set of existing approaches [2, 3, 4, 5] adopt an open-loop design that lacks mechanisms to verify the correctness of the operation code generated by LLMs.
+Such open-loop designs can leave questions such as cumulative error and temporal action consistency unsolved [15], and hence affect the reliability of their generated code, especially when handling complex tasks.
+Recently, closed-loop designs have been proposed in LLM-driven robotic operation code generation [6, 7, 8, 9, 10, 11], which enable evaluation of generated code and provide feedback for refinement, thereby steering robot execution towards its task objectives.
+Nevertheless, these approaches still face limitations from different aspects, including:
+  • Using numerical state observations for evaluation: approaches proposed in ref [6, 7] use numerical state observations in scenarios for evaluation and feedback generation.
+Due to the relatively limited numerical reasoning and processing capabilities of LLMs [16], an LLM-driven evaluation can overlook the latent semantics of robot actions and fail to achieve a comprehensive assessment of its movements and dynamics, leading to unreliable code refinement feedback.
+  • Physical execution for evaluation: closed-loop approaches, such as ref [8, 9, 7], employ physical execution of LLM-generated code to generate state observations for evaluation and code refinement.
+Although these approaches can be acceptable for robots that are not likely to cause physical damage during the refinement, their adoption in the context of UAV operation becomes risky because the physical execution of UAVs with code that potentially contains errors for refinement can easily lead to UAV crashes and safety concerns [17].
+  • Self-refinement using a single LLM: existing closed-loop approaches [10, 11] that a single LLM can be used for both code generation, evaluation, and refinement.
+However, self-evaluation may amplify its errors and lead to degradation in task success rate [18], which is also validated in the experimental evaluation of this work.
+To overcome these limitations, this paper proposes a novel closed-loop framework to further enhance the reliability of LLM-driven UAV operation code generation.
+Specifically, our framework adopts a duo-LLM design, in which one LLM is configured as the Code Generator and the other one is configured as the Evaluator.
+To compensate for LLM’s limitation in numerical reasoning, we propose a state transformation algorithm that converts numerical state observations into semantic observations by embedding UAV dynamics within the context of semantic trajectory observations during closed-loop evaluation.
+In addition, our framework enables simulation-based evaluation and refinement of LLM-generated UAV operation code before physical deployment.
+This feature enables efficient refinement for UAV operation code generation, and more importantly, eliminates the safety risks raised by potential incorrect code execution during refinement with physical UAV execution.
+Align with existing research [6, 7, 8, 9, 10, 11], this paper also focuses on using LLMs to generate operation code for deployment and task execution with the complete task description provided, but not real-time interactive UAV controls.
+The high-level design of our framework is illustrated in Fig.
+2.
+In our framework, the LLM configured as the Code Generator synthesizes or refines UAV operation code based on task description or evaluation feedback, and the LLM configured as the Evaluator assesses code performance and provides feedback to guide code refinement.
+The process unfolds as follows: 1) The Code Generator first produces an initial code based on the task description; 2) The code is then execute in simulation to produce raw numerical state observation, which are subsequently transformed into a semantic observations; 3) The Evaluator evaluates the semantic observations against the original task specification and returns evaluation feedback; 4) If the evaluation confirms that the observations has met the task objectives, the code will deploy on the UAV for task execution.
+Otherwise, the evaluation feedback is returned to the Code Generator for iterative refinement.
+To evaluate the effectiveness of our proposed framework, we conducted extensive experiments on UAV operation tasks with varying levels of complexity.
+The experimental results show that our framework achieves a $100\%$ success rate for basic tasks.
+For advanced tasks with a higher complexity, our framework achieves a $98.5\%$ completeness and an $85\%$ success rate, i.e., $98.5\%$ of required actions in all evaluated tasks are completed correctly and $85\%$ of tasks are entirely completed without any error.
+Compared with baseline approaches [5, 10, 7], our framework greatly improves both success rate and completeness when the task complexity increases.
 Figure 2: Illustration of LLM-driven closed-loop UAV Operation and Refinement.
-Inspired by control theory, more recent studies have adopted a closed-loop architecture that enables evaluation of generated code and provides feedback for refinement, thereby steering robot execution towards prescribed task objectives [20, 21, 22, 23, 24, 25, 26].
-Nevertheless, challenges are prone.
-First, existing approaches prompt LLM with numerical state observations in scenarios that require insights into robot dynamics for evaluation [24, 26].
-Given that LLMs exhibit limited numerical reasoning and processing capabilities [27], an LLM-driven evaluation may overlook the latent semantics of robot actions and fail to achieve a comprehensive assessment of the robot’s movements and dynamics, leading to unreliable code refinement feedback.
-Second, current closed-loop approaches require physical robot execution to generate state observations for evaluation [22, 24].
-Such an approach may be acceptable for ground robots where occasional errors rarely cause irreparable damage.
-However, it poses risks for aerial platforms, as even slight control deviations can precipitate UAV crashes.
-To address these challenges, we propose a novel closed-loop feedback and refinement framework that leverages semantic observations to enhance the reliability of LLM-driven UAV control as depicted in Fig.
-1.
-Our framework comprises two LLM modules: a Code Generator, which synthesizes or refines UAV operation code based on task description or evaluation feedback; and an Evaluator, which assesses code performance and provides feedback to guide code refinement.
-The process unfolds as follows: 1) The code generator first produces an initial code based on the task description; 2) The code is then execute in simulation to produce raw numerical state observation, which are subsequently transformed into an NL trajectory observation; 3) The evaluator compares NL trajectory observation against the original task specification and returns evaluation feedback; 4) If the evaluation confirms the trajectory meets the task objectives, the code will deploy on the UAV for mission execution.
-Otherwise, the evaluation feedback is returned to the code generator for iterative refinement.
-By introducing NL trajectory representations rather than numerical state information, our method significantly enhances the evaluator’s comprehension of the UAV’s state dynamics, thereby improving the reliability and precision of the evaluation feedback.
-Moreover, our method enables the refinement process in simulation, which eliminates the risks to UAVs caused by incorrect code execution during refinement, and hence supports safe and efficient development.
-We conduct extensive experiments on UAV operation tasks with varying levels of complexity to evaluate the effectiveness of our proposed framework.
-The experimental results show that our framework achieves a $100\%$ success rate and $100\%$ completeness for basic operations.
-For complex tasks, our framework achieves an $85\%$ success rate and $98.5\%$ completeness.
-Compared with baseline approaches [15, 25, 24], our framework greatly improves both success rate and completeness when the task complexity increases.
-The contributions of our work can be summarized as follows:
-  • Closed-Loop Refinement: A novel closed-loop feedback and refinement framework with a code generator and an evaluator designed to enhance the reliability of LLM-driven UAV control.
-  • Semantic Encoding: A novel algorithm for converting numerical state observations into step-wise NL trajectory descriptions, thereby enhancing the evaluative reasoning process.
-  • Safe Development: A simulation-based LLM-generated UAV code evaluation that replaces physical deployment to prevent potential UAV crashes in code refinement.
-  • Comprehensive Evaluation: A comprehensive experimental evaluation that demonstrates the performance of our framework with a detailed analysis of design factors.
 The remainder of this paper is structured as follows: Section II reviews and discusses existing research.
 Section III presents our problem formulation, which is followed by the detailed design of our proposed framework in Section IV.
-Section IV-D provides the evaluation and analysis of our framework.
-This paper is concluded in Section VI.
+Section V provides experimental evaluation and analysis of our framework.
+We discuss the applicability of our framework and future work in Section VI, and conclude this paper in Section VII.
 II Related Work
 II-A LLM for Robotics
-LLMs have demonstrated exceptional capabilities in semantic understanding and context generation, making them invaluable in robotics.
-ChatGPT for Robotics [11] and Code as Policies [18] have shown significant progress in LLM to control various types of robots.
-In the domain of aerial platforms, TypeFly [12] introduces an end-to-end LLM framework for piloting drones via human languages, while REAL [28] integrates LLMs into a mission planning and control framework.
-Moreover, continual advancements in LLMs’ reasoning and code generation capabilities [5, 29] are opening new opportunities for a variety of LLM-driven robotic tasks.
-Despite these advances, robotic performance continues to be limited by the reasoning capacity of LLMs, which frequently demonstrate unreliability when faced with tasks requiring complex reasoning [11].
+Continual advancements in LLMs [19, 20] have demonstrated their exceptional capabilities in semantic understanding and context generation, making them promising when integrating with various robotics tasks.
+Research works from Microsoft [2] and Google [21] demonstrated how publicly accessible LLMs can be leveraged to generate operation code for various robotic tasks, including UAVs.
+Since then, a set of research works [3, 4, 5, 6, 7, 8, 9, 10, 11] have been proposed to enhance the quality and reliability of LLM-driven robotic tasks, especially with the integration of prompt engineering techniques [13] that help steer LLMs’ behavior and align their outputs with the tasks’ objectives.
+In ref [3], an end-to-end LLM framework is proposed that enables semantic planning with applications in quadcopter drones.
+This paper also introduced a specific programming language (MiniSpec) to improve the code generation efficiency of LLMs.
+In ref [4], a prompt-engineered robot control solution is proposed that synthesizes input from both LLMs and humans.
+This work aims to have the robotic system behavior programmed by emphasizing prompting LLMs and avoiding manipulating code.
+Recently, a prompt framework with enhanced reasoning, named GSCE [5] was proposed to better support LLM-driven UAV control.
+GSCE integrates domain-specific guidelines, constraints, and examples into its system prompts when configuring the LLM, and hence improves its reliability of constraint-compliant code generation.
+While these methods all improve the LLM-driven code generation for robotic tasks from different aspects, they all adopt an open-loop design framework without error correction and adjustment for the generated code.
+Such a design can lead to potential cumulative errors and temporal action consistency issues [15] in the code generation.
+To address the concerns from the open-loop design, different closed-loop refinement strategies are increasingly adopted by recent research towards LLM-driven code generation for robotic tasks.
+In ref [6, 7], numerical observation of the robot’s execution status, such as actuator reading, sensor reading, and system information (e.g., location), is utilized to construct the feedback for LLMs.
+These numerical observations-based feedback can positively contribute to the quality of LLM-generated operation code, however, their contribution is restricted when trying to assess the latent semantics of robot actions for a comprehensive assessment.
+This is because LLMs have relatively limited numerical reasoning and processing capabilities [16].
+In addition, to obtain accurate observations and leverage them for the refinement of LLM-generated code, executing the code with physical experiments is employed for feasible robotic tasks.
+For example, CoPAL [8] employs physical error feedback and run-time error feedback to refine the code when errors occur during physical execution.
+In ref [9], precondition errors are integrated into a prompt that aims to repair plans when the robot fails skill execution.
+While evaluating code with physical experiments can provide accurate observations of errors and misalignment between LLM-generated code and task objectives that enable the construction of an effective refinement plan, it does not fit when moving to the context of many UAV tasks.
+Physical execution of code that may contain errors, especially before any refinement, has a high chance of causing the UAV to crash [17].
+This not only increases the hardware cost for the experiment but also can lead to safety concerns.
+Therefore, our proposed framework enables the execution of LLM-generated UAV operation code in the simulation environment to resolve the safety concerns in the refinement process.
+Moreover, the state transformation design in our framework enables semantic observations of the execution rather than numerical-only observation, and hence empowers LLMs in the code evaluation and feedback generation for refinement.
+In recent research, other strategies to further improve the performance of closed-loop design have also been explored.
+For example, ref [22] utilizes human feedback to revise the code when misalignment in the generated plan is detected.
+Although human feedback provides high-quality input for refinement, it limits the automation of the system and the performance depends on the expertise of the user.
+In addition, existing research [6, 7, 8, 9] separates the configuration of code generator and evaluator on different LLMs.
+This design aims to minimize the amplification of errors and bias [18] that can be caused by using a single LLM for both code generation and evaluation in previous research [10, 11].
+In our framework, we also use two separate LLMs and have customized configurations to improve their capabilities as the code generator and evaluator, respectively.
 II-B Prompt Engineering
@@
 Prior studies have embedded CoT within the examples to guide LLMs through reasoning step-by-step via in-context learning [26, 25].
-In this study, we integrate both CoT and in-context learning strategies to enhance the quality of code generation and evaluation in LLM-driven UAV control.
-II-C LLM-based Feedback, Evaluation, and Refinement
-ChatGPT for Robotics emphasizes the necessity of human oversight for code generation [11], reflecting the fact that robotic performance remains constrained by LLMs’ reasoning limitations.
-Moreover, LLMs exhibit sensitivity to input perturbations [33, 34, 35].
-Thus, recent research efforts adopt closed-loop refinement strategies to verify the LLM generation for improving the reliability of LLM-driven control.
-For instance, PluginSafety [36] designs a module that enforces predefined logical constraints evaluation to ensure LLM-generated content adheres to predefined constraints.
-More recent studies employ corrective feedback to iteratively refine generated code.
-Some integrate human or automated checker feedback [20, 21, 37, 38] to re-prompt the LLM and improve reasoning accuracy.
-While others monitor the robot status during execution and provide feedback, such as runtime errors, robot constraints violations [39, 22, 23, 24].
-The evaluation processes generally fall into two categories: self-evaluation and external evaluation.
-SELF-REFINE [25, 39] leverages a single LLM to assess and refine its own outputs based on feedback, whereas external evaluation approaches employ a secondary LLM or external evaluator to generate corrective feedback [36, 22, 23, 24].
-In our approach, numerical state information is first transformed into NL observations, and then we use an external LLM-evaluator to provide feedback for iterative refinement.
-Figure 2: An example of how our LLM-driven closed-loop feedback and refinement framework works during solving a task that requires examining a square area.
-In the first loop, the evaluator identifies deviations in the NL trajectory observation (action 12), providing error feedback into the code generator to refine the code.
-In the second loop, the evaluator confirms refined code matches the task, and the code is deployed on the UAV for mission execution.
+In this study, we integrate both CoT and in-context learning strategies to enhance the quality of code generation and evaluation.
 III Problem Formulation
@@
 The state transitions of $l$ yield state observations $O$, and evaluation $E=\langle O,P\rangle$ assesses whether $O$ satisfies the problem objectives and outputs $F$ as feedback.
-If $F$ is ”YES”, the sequence of actions $l$ is deemed a valid solution solution that flies the UAV from $s_{0}$ to the goal state $(s_{n}\equiv G$); otherwise, $l$ is iteratively refined based on the feedback $F$.
+If $F$ is “YES”, the sequence of actions $l$ is deemed a valid solution solution that flies the UAV from $s_{0}$ to the goal state $(s_{n}\equiv G$); otherwise, $l$ is iteratively refined based on the feedback $F$.
+Figure 3: Overall design of LLM-driven closed-loop UAV operation code generation with semantic observations.
+An example task of examining a square area is presented.
+In the first loop, the evaluator identifies deviations in the semantic trajectory observation (action 12), providing error feedback into the code generator to refine the code.
+In the second loop, the evaluator confirms refined code matches the task, and the code is deployed on the UAV for mission execution.
 IV Methodology
 IV-A Overview
-The overall idea of our proposed reliable LLM-driven closed-loop control framework for UAV operations is illustrated in Fig.
-2 and Algorithm 1.
-Given a UAV operation task description by the user, the LLM configured as the code generator with enhanced reasoning capabilities towards UAV operation first produces the initial version of the operation code.
-A simulation-based execution of the code is then performed to obtain the numerical state observations.
-To close the gap between numerical state observations and the LLM-driven evaluation, a state transformation is designed that transforms numerical state observations into an NL trajectory observation.
-The evaluator analyzes the trajectory observation together with the task description.
-If the trajectory satisfies all task objectives in evaluation, the operation code is ready for deployment; otherwise, the evaluator provides feedback to the code generator for refinement.
-This process iterates until the task objectives are achieved or the maximum number of iterations is reached.
-We now provide the detailed design of our code generation, trajectory observation, evaluation, feedback, and refinement in the following sections.
-Input: $Task$: Task description
-Output: $Code$: Code script
-$Code\leftarrow$ Code Generator$(Task)$
+The overall design of our proposed reliable LLM-driven closed-loop control framework for UAV operations is presented in Fig.
+3, in which an example task of examining a square area is also added.
+In addition, we also summarize the closed-loop feedback and refinement process of our LLM-driven code generation in Algorithm 1.
+Specifically, given a UAV operation task description ($P$) by the user, an LLM configured as the code generator with enhanced reasoning capabilities towards UAV operation first produces the initial version of the operation code $l$.
+A simulation-based execution of the code $l$ is then performed to obtain the numerical state observations.
+To close the gap between numerical state observations and the LLM-driven evaluation, a state transformation is designed that transforms numerical state observations into a semantic trajectory observation $O$.
+After that, the evaluator analyzes the semantic observation together with the task description and output evaluation feedback $F$ that depict the deviations from the $P$ in $O$ (if any).
+If the trajectory $O$ satisfies all task objectives in evaluation, the operation code is ready for deployment; otherwise, the code generator re-generates the refined $l$ according to the evaluator’s feedback $F$.
+This “generation-simulation-evaluation” process iterates until the evaluation confirms task objectives are achieved or the maximum number of iterations is reached.
+We now provide the detailed design of our code generation, trajectory observation, evaluation, as well as feedback and refinement in the following sections.
+Input: $P$: Task description
+Output: $l$: Code script
+$l\leftarrow$ Code Generator(P)(P)
 $itr\leftarrow 1$
 while $itr\leq$ maximum round do
-       $O\leftarrow$ State Transformation$(Code)$
-      $F\leftarrow$ Evaluator$(O,Task)$
+ $O\leftarrow$ State Transformation(l)(l)
+ $F\leftarrow$ Evaluator(O,P)(O,P)
  if “YES” in $F$ then
@@
  else
-             $Code\leftarrow$ Code Generator$(F)$
+    $l\leftarrow$ Code Generator(F)(F)
    end if
@@
 end while
-return $Code$
+return $l$
 Algorithm 1 Closed-loop Feedback and Refinement
@@
 These observations are numerical representations of the UAV’s states, including the UAV’s position and orientation.
-However, numerical observations offer limited semantic insight for an LLM-based evaluator because LLM lacks prior knowledge about interpreting the semantic meaning of numerical state tuples [27].
-To bridge this gap, we transform each numerical state observation into a step-wise NL description and aggregate these into an NL trajectory observation.
-As shown in Fig.
-1, NL description captures the semantic meaning of the state changes, such as “Move 5 meters West while facing North, the UAV moves to [0, 5]”.
-This transformation provides the evaluator with a clear understanding of the semantic significance of each action and overall trajectory dynamics, thereby improving the evaluator’s ability to identify deviations and generate precise feedback for code refinement.
-Given the LLM-generated UAV operation code executed in simulation, Algorithm 2 presents the transformation from numerical state observations to the NL trajectory observation $O$.
-For each action identified in the $Code$, its execution in the simulation will be observed to obtain the corresponding numerical UAV state observations.
-Specifically, for each action $a$, the simulator first executes $a$, if a runtime error occurs, the algorithm collect and log errors alongside the corresponding action identifier in $O$; otherwise, the algorithm computes the difference between the current state $s_{\mathrm{current}}$ and the last state $s_{\mathrm{last}}$, denoted as $\Delta s=s_{\mathrm{current}}-s_{\mathrm{last}}$.
-For each element in the state difference $\Delta s$, the algorithm appends the corresponding NL description of orientation and changes in yaw, x-axis, y-axis, or z-axis to $O$.
-The algorithm also appends the UAV’s NED position to $O$.
-Upon completion of all actions, the algorithm returns the complete NL trajectory observation $O$.
-Input: $Code$: Code script
-Output: $O$: Trajectory observation
+However, numerical observations offer limited semantic insight for an LLM-based evaluator because LLM lacks prior knowledge about interpreting the semantic meaning of the numerical state vector [16].
+To bridge this gap, we transform each numerical state observation into a step-wise semantic description and aggregate these into a semantic trajectory observation.
+As an example shown in Fig.
+4, the semantic description captures the semantic meaning of the state changes and UAV position, such as “Move 5 meters East while facing North, the UAV moves to [0, 5, -11.5]”.
+This transformation provides the evaluator with a clear understanding of the semantic significance of each action and overall trajectory dynamics, thereby improving the evaluator’s ability to identify deviations and generate precise feedback for guiding the code generator on code refinement.
+Algorithm 2 presents the transformation from numerical state observations to the semantic trajectory observation $O$.
+For each action identified in the $l$, its execution in the simulation will be observed to obtain the corresponding numerical UAV state observations $x,y,z,\theta]$ and then transformed to semantic trajectory descriptions as the example illustrated in Fig.
+4.
+Specifically, for each action $a$, the simulator first executes $a$, if a runtime error occurs, the algorithm collect and log errors alongside the corresponding action identifier in $O$; otherwise, the algorithm computes the state transition $T$ between the current state $s_{\mathrm{current}}$ and the last state $s_{\mathrm{last}}$, denoted as $T=s_{\mathrm{current}}-s_{\mathrm{last}}$.
+For each element in the state transition $T$, the algorithm appends the corresponding semantic description of orientation and changes in yaw, x-axis, y-axis, or z-axis to $O$.
+The algorithm also appends the UAV’s NED position vector to $O$ for embedding the positional information.
+Upon completion of all actions of $l$, the algorithm returns the complete semantic trajectory observation $O$.
+Figure 4: An example of transforming numerical state observations into semantic trajectory observation. UAV state vector of $x,y,z,\theta]$ denotes (North, East, Down, Yaw).
+Input: $l$: Code script
+Output: $O$: Semantic trajectory observation
 $O\leftarrow\emptyset$
-foreach  $action$ in $Code$  do
-       $s_{\mathrm{last}}\leftarrow$ State Observation $(x,y,z,\theta)$
-      Execute $action$
-      if error executing $action$ then
-             $O\leftarrow O$ + “Error in executing $action$ with error message $error$
+foreach  actionaction in $l$ do
+ $s_{\mathrm{last}}\leftarrow$ Numerical state observation $x,y,z,\theta]$
+  Execute actionaction
+ if error executing actionaction then
+    $O\leftarrow O$ + “Error in executing actionaction with error message errorerror
    end if
-      $s_{\mathrm{current}}\leftarrow$ State Observation $(x,y,z,\theta)$
-      $\Delta s\leftarrow s_{\mathrm{current}}-s_{\mathrm{last}}$
-      foreach $element$ in $\Delta s$ do
-             if $element$ in orientation then
-                   $O\leftarrow O$ + Rotate $element$ degrees in Yaw.
+ $s_{\mathrm{current}}\leftarrow$ Numerical state observation $x^{\prime},y^{\prime},z^{\prime},\theta^{\prime}]$
+ $T\leftarrow s_{\mathrm{current}}-s_{\mathrm{last}}$
+ foreach elementelement in $T$ do
+    if elementelement in orientation then
+       $O\leftarrow O$ + Rotate elementelement degrees in Yaw.
     else
-                  if $element$ in NED axis then
-                         $O\leftarrow O$ + “Move $element$ meters in North, East, or Down while facing $s_{\mathrm{current}}[\theta]$.”
+       if elementelement in NED axis then
+          $O\leftarrow O$ + “Move elementelement meters in North, East, or Down while facing $s_{\mathrm{current}}[\theta]$.”
          end if
@@
 IV-D Evaluation
-In our framework, we design the evaluation by leveraging the LLM’s semantic reasoning capabilities to analyze the NL trajectory observation.
-The evaluator aims to identify deviations (if any) between the NL trajectory observation and the task description, and then construct feedback based on the deviations.
+In our framework, we design the evaluation by leveraging the LLM’s semantic reasoning capabilities to analyze the semantic trajectory observation.
+The evaluator aims to identify deviations (if any) between the semantic trajectory observation and the task description, and then construct feedback based on the deviations.
 Instead of using the same LLM used for code generation for self-feedback evaluation, our design employs a separate LLM to prevent the same LLM from amplifying its bias [18].
-To improve the precision of LLM’s evaluation, we adopt the design principles of the code generator by configuring the evaluator as an LLM agent with a structured system prompt.
+To improve the accuracy of LLM’s evaluation, we adopt the design principles of the code generator by configuring the evaluator as an LLM agent with a structured system prompt.
 However, unlike the code generator, the evaluator does not produce code but instead generates evaluation feedback.
 To address this issue, we construct the evaluator’s system prompt with roles, rules, and references, as depicted in Fig.
-3.
-Specifically, the roles define the LLM agent as an evaluator that compares the NL trajectory observation against the task description, while the rules serve as explicit evaluation criteria to direct the LLM when reasoning about trajectories.
-Additionally, the references are exemplar trajectories that serve as benchmarks to support the evaluation of NL trajectory observations.
+5.
+Specifically, the roles define the LLM agent as an evaluator that compares the semantic trajectory observation against the task description, while the rules serve as explicit evaluation criteria to direct the LLM when reasoning about trajectories.
+Additionally, the references are example trajectories that serve as benchmarks to support the evaluation of semantic trajectory observations.
 Furthermore, the system prompt requires the evaluation outcome to be informative and constructive by providing explanations about the depicted deviations.
 As demonstrated in Fig.
-2, given a NL trajectory observation and a task description, the structured system prompt enables the LLM-evaluator to identify deviations from the task description and explain the errors in the trajectory observation (e.g., “Action 12 flies West instead of East”).
+3, given a semantic trajectory observation and a task description, the structured system prompt enables the LLM-evaluator to identify deviations from the task description and explain the errors in the trajectory observation (e.g., “Action 12 flies West instead of East”).
 Consequently, the code generator is aware of the errors and could correctly refine the code in the subsequent refinement process.
 Figure 5: The evaluation process begins by configuring the LLM agent (evaluator) via a system prompt framework that includes roles, rules, and references.
-Next, the trajectory observations and the task description are provided to the evaluator for comparison.
+Next, the semantic observations and the task description are provided to the evaluator for comparison.
 IV-E Feedback and Refinement
@@
 They leverage the two LLM agents designed in section IV-B and section IV-D and form a closed-loop as illustrated in Fig.
-2.
+3.
 Specifically, the code generator interprets the evaluation feedback $F$ and refines the generated code accordingly.
-Then, the refined code is executed in the simulation again to produce a new NL trajectory observation $O^{\prime}$ for re-evaluation.
-The feedback and refinement loop continues iteratively until the evaluator outputs ”YES”, indicating that $E$ confirms that the refined action sequence $l$ successfully solves the task problem $P$ and the UAV operation code is ready for deployment.
+Then, the refined code is executed in the simulation again to produce a new semantic trajectory observation $O^{\prime}$ for re-evaluation.
+The feedback and refinement loop continues iteratively until the evaluator outputs “YES”, indicating that $E$ confirms that the refined action sequence $l$ successfully solves the task problem $P$ and the UAV operation code is ready for deployment.
 In addition, if the maximum number of iterations is reached without a successful evaluation, the evaluator suggests human intervention.
-V Experiment
-The goal of this experiment is to validate the effectiveness of our proposed LLM-driven closed-loop framework in enhancing the reliability of LLM-driven UAV operation code generation.
-To achieve this, the experiment is designed with three primary goals: (1) to evaluate the performance of our closed-loop feedback and refinement framework in LLM-driven UAV tasks; (2) to analyze the impact of different design choices on the performance of our framework; (3) to assess the behavior of our framework under diverse task scenarios.
+V Experiment and Evaluation
 V-A Experimental Setup
-V-A1 Environment
-To ensure safe code execution without risking real-world failures, we conduct experiments in the “AirSim” simulator [41], which realistically simulates UAV (e.g., a quadcopter) dynamics and environmental interactions.
-During the experiment, UAV state information was fully accessible from the simulator in real time and utilized for closed-loop feedback and evaluation purposes.
-V-A2 Tasks
-We propose Basic and Advanced task sets to rigorously evaluate the performance of LLM-driven control code generation across varying levels of operational complexity [^1] The complete list of tasks is available at https://github.com/ai-uavsec/CLGSCE.
-The Basic set comprises 44 tasks adopted from GSCE [15].
-Each task involves one or two intuitive discrete movements to evaluate LLM’s few-stage planning and basic spatial reasoning (e.g., “Turn 90° clockwise, then fly 4 meters left in the drone’s body frame”).
-In contrast, the Advanced set consists of 20 tasks that emulate real-world UAV deployment (e.g., “Take off and fly up 5 meters.
+In our experiment, the implementation of our proposed framework and comparison methods [5, 10, 7] uses the OpenAI “o3-mini” (o3-mini-2025-01-31) model [20] as the foundational LLM.
+Based on the “o3-mini”, two separate LLMs are configured as the code generator and evaluator in our framework.
+The widely used UAV simulator, “AirSim” [27], is tailored to support the simulation component of our framework for code execution and state observations.
+The UAV model selected for experiment and evaluation is the default quadcopter in AirSim.
+The detailed implementation of our framework and experiment is available at our project website  [^1] Project Website: https://github.com/ai-uavsec/CLGSCE.
+UAV Task Dataset: In our experiment, two categories of UAV tasks are selected to evaluate the performance of our framework under different operational complexities, i.e., Basic tasks and Advanced tasks.
+Each Basic task involves one or two intuitive actions to evaluate LLM’s few-stage planning and basic spatial reasoning (e.g., “Turn 90° clockwise, then fly 4 meters left in the drone’s body frame”).
+We use the Basic task dataset established by existing research [5], which contains 44 tasks in total.
+Differently, the Advanced tasks in our experiment require complex reasoning about the UAV’s position and orientation dynamics and typically require 6-19 actions to reach the goal state.
+We established the Advanced task dataset by emulating real-world UAV operation tasks, which consists of 20 tasks with different complexities.
+As summarized in Table I, there are four levels of complexity: “Low”, “Medium”, “High”, and “Very High”.
+The “Low” tasks involve flying simple geometric patterns.
+The “Medium” tasks introduce orientation requirements while flying the patterns (e.g., always orienting to the flying direction).
+The “High” tasks require flying complex patterns with orientation requirements (e.g., fly a figure 8 while orienting in the flying direction).
+The “Very High” increases the complexity by adding requirements to the patterns (e.g., inspect an area twice with the second inspection flies the path reversely).
+The complete list of both Basic and Advanced tasks is available at our project website \@footnotemark All tasks were manually validated in our experiment to avoid potential simulator-induced errors that could affect the evaluation process.
+TABLE I: Examples of Tasks in Advanced Task Dataset
+Complexity Num. of Tasks Avg. Steps Example Tasks
+Low 3 9 “You are going to scan square areas. First, take off and climb up 5 m. Fly a 5 m square (forward/right/back/left). Then shift 5 m south, and fly the same square pattern again.”
+Medium 9 11 “You are going to view the borders of square areas.
+Please take off and fly up 5 meters.
+You should fly in a square pattern with 5-meter sides by moving forward, right, backward, and left in world axis.
+Make sure the drone is oriented to the flying direction.”
+High 4 17.5 “You are going to fly a figure of 8.
+Please take off and fly up 5 meters.
+Fly a figure of 8 on a flat, horizontal plane with each side of 5 meters.
+The left square is on your left-rear side, and the right square is on your right-front side.
+You should begin with the left square by flying left.
+The right square should start from moving forward.
+Make sure the drone is oriented in the flying direction.”
+Very High 4 17.5 “You are going to examine a cube.
+Take off and fly up 5 meters.
+You will examine a square area.
 You should fly in a square pattern with 5-meter sides by moving forward, right, backward, and left.
-Make sure the drone is oriented to the flying direction.”).
-These tasks demand reasoning about the objectives in the task description and require 6-19 UAV actions to reach the goal state.
-The advanced task set is intended to evaluate LLM’s multi-stage planning and complex spatial reasoning.
-Additionally, all tasks in both task sets were manually validated to avoid potential simulator-induced errors that could affect the evaluation process.
-V-A3 Compared Methods
-Three recently proposed methods [15, 25, 24] are selected as the baselines for comparison with our approach, including
-  • GSCE: An open-loop approach in which the GSCE code generator synthesizes control code based on the task prompt, without any refinement. [15]
-  • LLM Self-feedback (Self-Refine): A single LLM performs both code generation and self-refinement.
-The LLM analyzes NL trajectory observations to detect deviations from the task and refines code based on its self-assessment.
-[25]
-  • Feedback with Numerical State Observations (Numerical): This method introduces a secondary LLM serving as evaluator designed according to section IV-D.
+To examine the square area, the drone should orientate perpendicular to the moving direction on each side of the square.
+Next, ascend another 5 meters and fly the square pattern in reverse order to examine the same area.”
+Baseline Methods for Comparison: Three recently proposed methods [5, 10, 7] are selected as the baselines for comparison with our proposed approach, including
+  • GSCE [5]: An open-loop approach in which the GSCE code generator synthesizes control code based on the task prompt, without any refinement.
+  • LLM Self-feedback (Self-Refine) [10]: A single LLM performs both code generation and self-refinement.
+The LLM analyzes semantic trajectory observations to detect deviations from the task and refines code based on its self-assessment.
+  • Feedback with Numerical State Observations (Numerical) [7]: This method introduces a secondary LLM serving as evaluator designed according to section IV-D.
 The evaluator detects deviations from numerical state observations and provides corrective guidance for the code generator for refinement.
-[24]
-All methods described above employ “o3-mini” (o3-mini-2025-01-31) [29] as the foundational LLM.
-The generated code was then deployed on a quadcopter in AirSim [41] to evaluate the performance of each method.
-To mitigate the randomness of LLM’s output [42], we repeat each task three times in the experiment.
-V-B Evaluation Setup
-V-B1 Metrics
-The performance is evaluated using two metrics as GSCE [15]: Completeness and Success Rate (SR).
-Completeness quantifies the proportion of goal states successfully reached, calculated as the ratio of the total count of correct actions to the total number of required actions of the ground truth trajectory in each task.
-The completeness offering insights into performance across the intermediate control process.
-$Completeness=\frac{\sum\text{correct actions}}{\mathrm{len}(l)}$ (1)
-SR is grounded in real-world operational requirements, wherein the UAV must fly the trajectory precisely as humans instructed.
-As formalized in Eq.
-2, a task is counted as a success only when the UAV reaches the final state while following the correct sequence of actions that lead to the ground truth trajectory.
-$Success=\begin{cases}1,&Completeness=1,\\ 0,&Completeness<1.\end{cases}$ (2)
-V-B2 Ground Truth
-The ground truth is represented by a list of state transitions.
-It is a vector of four elements: $[x,y,z,\theta]$, where $x$, $y$, and $z$ denote the UAV’s position changes in the North, East, and Down axes, and $\theta$ represents yaw rotation.
+V-B Evaluation Metrics
+In our evaluation, two key metrics Completeness and Success Rate (SR) are used for performance measurement.
+Given a UAV task, the completeness measures the ratio of actions required by the task that are completed successfully, and the SR measures whether the entire task is completed successfully.
+The detailed definition of these two metrics, as well as the ground truth used in our evaluation, is provided below.
+Completeness: As formalized in Eq.
+1, completeness for task $i$ computed as the ratio between the count of actions correctly executed ($a^{\mathrm{correct}}_{i}$) and the total number of actions in the task’s ground truth action sequence ($l^{\mathrm{gt}}_{i}$).
+This metric offers insights into performance across the intermediate control process, where a higher completeness indicates a greater proportion of the task that has been successfully completed.
+$Task\_Completeness_{i}=\frac{|a^{\mathrm{correct}}_{i}|}{|l^{\mathrm{gt}}_{i}|}$ (1)
+Given $n$ tasks, the completeness is calculated as:
+$Completeness=\frac{1}{n}\sum_{i=1}^{n}Task\_Completeness_{i}$ (2)
+SR: As formalized in Eq.
+3, a task is counted as a success only when the UAV reaches the final state while following the correct sequence of actions that lead to the ground truth trajectory.
+Then we compute the SR across $n$ tasks as Eq.
+4.
+This metric is grounded in real-world operational requirements, wherein the UAV must perform the task precisely as instructed.
+$Success_{i}=\begin{cases}1,&Task\_Completeness_{i}=1,\\ 0,&Task\_Completeness_{i}<1.\end{cases}$ (3)
+$SR=\frac{1}{n}\sum_{i=1}^{n}Success_{i}$ (4)
+Ground Truth: The ground truth is represented by a list of state transitions.
+It is a vector of four elements: $x,y,z,\theta]$, where $x$, $y$, and $z$ denote the UAV’s position changes in the North, East, and Down axes, and $\theta$ represents yaw rotation.
 We assume the UAV’s state is fully observable at all times, thereby ensuring that each transition can be precisely monitored and compared against ground truth.
@@
 V-C1 Overall Result
-The average SR and completeness across Basic and Advanced tasks are summarized in Table I.
+The average SR and completeness across Basic and Advanced tasks are summarized in Table II.
+The results reflect the average value of three repetitions of experiments to mitigate the impact of potential randomness of LLM’s output [28].
+TABLE II: SR and Completeness of Proposed Methods
+Basic Task Advanced Task SR Completeness SR Completeness GSCE [5] $\% 100.0100.0\%$ $66.7\%$ $88.9\%$ Self-Refine [10] $\% 100.0100.0\%$ $50.0\%$ $74.3\%$ Numerical [7] $95.5\%$ $97.3\%$ $73.3\%$ $92.4\%$ Ours $\% 100.0100.0\%$ $\mathbf{85.0\%}$ $\mathbf{98.5\%}$
 For the Basic tasks, all methods showed remarkable performance.
-Specifically, the GSCE, Self-Refine, and Ours methods achieved $100\%$ in both SR and completeness, while the Numerical achieved $95.5\%$ SR and $97.3\%$ completeness.
-This demonstrates their ability to generate correct code for a few-stage planning and basic spatial reasoning.
-However, the slightly lower performance of the Numerical method suggests that relying on numerical state observations can occasionally lead to misinterpretation in evaluation, resulting in incomplete or incorrect refinements.
-Performance on the Advanced tasks declines notably across all methods as a result of increased task complexity.
-The open-loop GSCE method achieves only $66.7\%$ SR, indicating that refinement becomes critical as task complexity increases.
-The Self-Refine method performs worse, with $50.0\%$ SR and $74.3\%$ completeness, proving that relying on a single LLM for both generation and self-evaluation amplifies bias [40], resulting in introducing additional errors in refinement.
-The Numerical method achieves $73.3\%$ SR and $92.4\%$ completeness, showing that it can detect some deviations; however, its numeric perspective still fails to provide reliable evaluation feedback for refinement.
-In contrast, our proposed method substantially outperforms all baselines, achieving $85.0\%$ SR and $98.5\%$ completeness on Advanced tasks.
-The overall results demonstrate that our proposed framework significantly enhances the reliability of LLM-driven UAV operation code generation, particularly for complex tasks that require multi-stage planning and complex spatial reasoning.
-TABLE I: Average SR and Completeness of Proposed Methods
-Base Task Advanced Task SR Completeness SR Completeness GSCE [15] $100.0\%$ $100.0\%$ $66.7\%$ $88.9\%$ Self-Refine [25] $100.0\%$ $100.0\%$ $50.0\%$ $74.3\%$ Numerical [24] $95.5\%$ $97.3\%$ $73.3\%$ $92.4\%$ Ours $100.0\%$ $100.0\%$ $\mathbf{85.0\%}$ $\mathbf{98.5\%}$
-V-C2 Failure Cases Analysis
-We further analyze the failure cases (cases that $Success=0$) in the Advanced task set.
-The analysis focuses on three aspects: (1) the number of failures and average completeness for each failed task ID with the respective ground truth action length (Fig.
-4); (2) the average completeness of each failed task ID (Fig.
-5); (3) the amount of task that experienced failure and the average completeness of all failed tasks (Table II).
-The results in Fig.
-4 reveal a clear trend that all methods exhibit a notable increase in failures as task complexity increases (i.e., more likely to fail on tasks with longer action length).
-Specifically, the baseline GSCE method failed in $11$ out of $20$ tasks ($55\%$ of the Advanced set), with an average completeness of $77.4\%$.
-This indicates that the GSCE method can reason through a substantial portion of the task correctly, however, it still struggles to maintain full alignment with the objectives of the task.
-This result highlights the need for improvement in the completeness of failed tasks to reduce overall failure rates.
-The Self-Refine method exhibited the weakest performance, with failures on $17$ tasks ($85\%$ of Advanced tasks).
-Moreover, the failures in Self-Refine were the least complete among all methods, demonstrating that performing evaluation and refinement on a single LLM not only fails to correct errors but may also introduce more errors.
-The Numerical method failed on $9$ tasks ($45\%$ of Advanced tasks) and achieved an average completeness of $69.5\%$ only.
-These results indicate that numerical state observations are insufficient for reliable code refinement and may even lead to erroneous refinement.
-Figure 4: Failure analysis across the four methods on the Advanced task set.
-The x-axis represents the task IDs that are sorted in ascending order based on the length of the task’s action steps.
-And the y-axis presents two metrics for each task: (1) the count of failures observed over three repetitions (left y-axis); (2) the length of the ground truth action steps (right y-axis).
-Figure 5: Average completeness for each failed task on the Advanced task set over four methods. The smaller the value, the less complete the failed task is.
-In contrast, our method achieves the best performance, failures occurred only on $4$ tasks ($20\%$ of Advanced tasks), and the completeness of failed tasks arose at $91.6\%$.
-Moreover, the completeness in Fig.
-5 demonstrates that even when failures occur, our method produces code that remains closely aligned with task objectives.
-These findings underscore the effectiveness of our method in refining errors and improving the completeness of LLM-driven UAV operation, which contributes to an increased success rate in handling complex UAV operation tasks.
-TABLE II: Amount of Task ID Experienced Failure and Average Completeness for All Failed Tasks
-Task ID Amount $\downarrow$ Failed Tasks Completeness $\uparrow$
-GSCE [15] $11$ $77.4\%$
-Self-Refine [25] $17$ $48.5\%$
-Numerical [24] $9$ $69.5\%$
-Ours $\mathbf{4}$ $\mathbf{91.6\%}$
-V-C3 Number of Refinement Iterations
-[…]
-Figure 6: Curves of average SR (left) and completeness (right) over the number of refinement iterations for the NL method on Advanced tasks.
+It is notable that the Numerical method [7] has a slightly lower performance with a $95.5\%$ SR and a $97.3\%$ completeness.
+This indicates that relying on numerical state observations can lead to misinterpretation in evaluation, even when handling simple tasks.
+When moving to the Advanced tasks that have a higher complexity, our method still achieved a reliable performance with an $85\%$ SR and a $98.5\%$ completeness.
+This result shows that $98.5\%$ of required actions in all evaluated tasks are completed correctly, and $85\%$ of tasks are entirely completed without any error.
+As a comparison, the reliability of baseline methods drops notably when the task complexity increases due to the limitations caused by factors such as amplified bias, limited numerical reasoning capabilities of LLMs, and open-loop design without refinement.
+While the impact of these factors is relatively negligible when handling basic tasks, their limitations are amplified when the task complexity increases, and affect the logical reasoning capabilities of LLMs to generate operation code correctly.
+Figure 6: Failure counts of different methods for each advanced task at different complexity levels
+Figure 7: Completeness of each failed task of different methods at different complexity levels
+TABLE III: Failure Cases Completeness in Each Level of Complexity
+Low Medium High Very Overall
+High
+GSCE [5] $90.9\%$ $81.9\%$ $0.7\% 77$$6\% 77.$$7.4\%$
+Self-Refine [10] $21.6\%$ $41.1\%$ $67.4\%$ $57.8\%$ $48.5\%$
+Numerical [7] $62.6\%$ $67.8\%$ $77.2\%$ $69.5\%$
+Ours $\mathbf{90.0\%}$ $\mathbf{94.4\%}$ $\mathbf{90.4\%}$ $\mathbf{91.6\%}$
+Failure Case Analysis: To uncover the detailed performance of each method in its failed advanced tasks, we performed an in-depth failure case analysis.
+In particular, we aim to understand how many actions are correctly completed when a task fails.
+As summarized in Table III, our method achieves a high average completeness of $91.6\%$ for all failed tasks, i.e., the operation code generated by our method can still complete $91.6\%$ of the required actions in the task and remains highly aligned with the task objectives even when errors occur.
+Differently, the average completeness of the compared baseline methods ranges from $48.5\%$ to $77.4\%$ for their failed tasks, which indicates errors occur for a significant portion of the required actions in these tasks.
+Fig.
+6 further details the failure count on each failed task of all methods in our three rounds of experiment with task complexity and length of actions information.
+Our method has failures on only four tasks during three rounds of experiments.
+As a comparison, the baseline methods failed 11 (GSCE [5]), 17 (Self-Refine [10]), and 9 (Numerical [7]) tasks, respectively.
+In addition, Table III and Fig.
+7 also show that our method has a stable high completeness of $90\%+$ for all failed tasks at different complexity levels.
+In contrast, the baseline methods have not only significantly lower average completeness ($21\%$ to $90.9\%$) on failed tasks at different levels but also a large variation in performance of different tasks.
+Based on our evaluation results, our method achieves a higher reliability compared with the baseline methods, not only because of its high confidence in generating fully correct operation code, but also its high completeness of the required actions of tasks when a small amount of errors occur.
+TABLE IV: Evaluator Configurations Description and Example System Prompts
+Evaluator Configuration Description
+(1) Roles This configuration defines the role for the LLM as the code evaluator, and its goal is to evaluate whether the UAV flight trajectory is aligned with the task description and output standardized feedback for the code generator.
+(2) Roles + Rules Besides the roles in (1), this configuration also includes specific rules for the LLM to follow.
+The rules introduce explicit evaluation criteria on the UAV’s actions, spatial positions, flight paths, coordinate system, and offset tolerance that direct the LLM’s evaluation of the UAV trajectory.
+(3) Roles + References Besides the roles in (1), this configuration also provides reference examples to support evaluation and generate feedback.
+These references include pairs of task descriptions and corresponding semantic trajectory observations.
+(4) Roles + Rules + References This configuration combines roles, rules, and references to provide comprehensive support information for the LLM to perform its job correctly.
+Example System Prompts
+Roles 1. You will reason about the difference between the task description and drone actions.
+2.
+If the task description and drone actions are aligned, you should output “YES” only.
+Otherwise, you should output “NO”; please output where the drone actions mismatch the task description if your answer is “NO”.
+Rules 1.
+For tasks that require perpendicularly examining a square area, you must check whether the drone is oriented perpendicularly inwards the square interior; facing toward the exterior is not allowed.
+2. You should reason about the spatial locations of square patterns from actions and ensure the drone flies the square pattern in the correct spatial position as outlined in the task description.
+3. The drone uses the NED coordinate system (world coordinates) represented as [x, y, z], the positive X axis is North/forward, the positive Y axis is East/right, and the negative Z axis is up/climb.
+References Here is a correct path of a 4-meter square pattern (forward/right/backward/left) for reference: 1.
+Fly 4 meters North in the world axis while facing north.
+The drone moves to [4, 0].
+2.
+Fly 4 meters East in the world axis while facing north.
+The drone moves to [4, 4].
+3.
+Fly 4 meters South in the world axis while facing north.
+The drone moves to [0, 4].
+4.
+Fly 4 meters West in the world axis while facing north.
+The drone moves to [0, 0].
+*Note: more examples of system prompts for roles, rules, and references are available at the project website: https://github.com/ai-uavsec/CLGSCE.
+V-C2 Number of Refinement Iterations
+In our proposed framework, the operation code generated by the Code Generator LLM will be refined through multiple iterations.
+The increased rounds of iterations provide additional opportunities to improve the code quality, however, they also introduce additional system cost (e.g., LLM API cost and processing time).
+To balance this tradeoff and identify an appropriate number of iterations, we performed experiments of our framework on advanced tasks with 0 to $10$ rounds of iterations.
+Our results presented in Fig.
+8 show the changes of SR and completeness with the different number of iterations.
+Specifically, both SR and completeness are boosted after the iteration is introduced for refinement, and then have a slowly increasing trend until the $6\textsuperscript{th}$ iteration.
+After that, the performance of the subsequent iterations becomes relatively stable in terms of both SR and completeness, and even declines slightly.
+This is because LLMs can exhibit sycophantic behavior and tend to adapt or change outputs to appear agreeable [29, 30], and thus may perform unnecessary changes after sufficient refinement has been performed.
+Our detailed examination identified that such a situation occurred on task 7 in our experiment.
+This observation demonstrates that iterative refinement can effectively contribute to the LLM-driven UAV operation code generation.
+Meanwhile, it is also necessary to impose an upper bound on refinement iterations to balance the tradeoff.
+According to our evaluation results, we set the maximum number of iterations to $6$.
+However, this number may need to be adjusted for different types of tasks.
+For example, setting a lower number when the tasks to complete have a relatively low complexity.
+Figure 8: Curves of SR (left) and completeness (right) over the number of refinement iterations for our method on advanced tasks.
 V-D Evaluator Configuration Analysis
-As the precision of the evaluation feedback directly influences the quality of code refinement, we also analyzed how different evaluator configurations affect evaluation precision.
-The evaluation precision is defined as the ratio of correct evaluations (i.e., correctly identifying whether the trajectory aligns with the task and providing a detailed explanation if deviations are identified) to the total number of evaluations conducted.
-To perform this analysis, we built a dataset from each Advanced task with one correct NL trajectory observation (Correct O) and one incorrect NL trajectory observation (Incorrect O).
-As presented in Table III, we tested four configurations of the evaluator system prompt: (1) roles; (2) roles $+$ rules; (3) roles $+$ references; (4) roles $+$ rules $+$ references.
-Moreover, each configuration was evaluated three times to reduce the randomness of LLM’s output [42].
-TABLE III: Evaluator Configurations
-Configuration (1) (2) (3) (4)
+As the accuracy of the evaluation feedback directly influences the quality of code refinement, we analyze how the evaluator’s design influences both evaluation accuracy and overall system performance.
+We evaluated four configurations of the evaluator’s system prompt, including (1) roles, (2) roles ++ rules, (3) roles ++ references, and (4) roles ++ rules ++ references.
+As specified in Section IV-D, the roles define the LLM to be an evaluator and serve as the essential component to ensure the LLM will perform the evaluation task.
+Thus, we include the roles in each configuration of the evaluator.
+For each configuration, we instantiate the evaluator’s system prompt by including only the indicated components.
+The detailed description of each evaluator configuration with sample system prompts for roles, rules, and references is provided in Table IV.
+TABLE V: Performance over Evaluator Configurations
+(1) (2) (3) (4)
 Roles

### v5

@@
 The experimental results show that our framework can achieve reliable UAV operations using LLMs, which significantly outperforms baseline methods in terms of success rate and completeness with the increase of task complexity.
-Index Terms:
-UAV, Large Language Model, Robotics, IoT.
 I Introduction
-Uavs, or drones, are increasingly being integrated into IoT ecosystems for various applications, such as surveillance and monitoring, smart cities, and infrastructure inspection [1].
+Uavs, or drones, are increasingly being integrated into IoT ecosystems for various applications, such as surveillance and monitoring, smart cities, and infrastructure inspection [uav_iot_applications].
 Their high mobility and rich sensing capabilities make them an ideal candidate for data collection, communication, and service delivery tasks in these applications.
@@
 Such a process requires substantial domain knowledge and can also be time-consuming.
-With LLMs’ exceptional semantic understanding and context generation capabilities, recent research has demonstrated the potential of leveraging LLM to facilitate the operation code generation process for a given robotic task [2, 3, 4, 5, 6, 7, 8, 9, 10, 11].
+With LLMs’ exceptional semantic understanding and context generation capabilities, recent research has demonstrated the potential of leveraging LLM to facilitate the operation code generation process for a given robotic task [ChatGPTRobotics, TypeFly, CLEAR, GSCE, incoro, InteractivePlanning, CoPAL, CAPE, SelfRefine, trust].
 As illustrated in Fig.
@@
 Figure 1: Pipelines of robotic operation code generation: (a) without LLMs and (b) with LLMs.
-Despite the benefits brought by LLM-driven UAV operation code generation, it also raises concerns about the reliability of LLM-generated code due to the challenges faced by LLM in terms of logical reasoning and complex decision-making [12].
+Despite the benefits brought by LLM-driven UAV operation code generation, it also raises concerns about the reliability of LLM-generated code due to the challenges faced by LLM in terms of logical reasoning and complex decision-making [UAVsmeetllmsoverviews].
 Different from text generation applications, where semantic-level accuracy is sufficient, the generation of UAV operation code requires the precise execution of sequential actions.
 Erroneous or invalid operation code can lead to unsafe UAV behavior and pose safety and security risks to the public, such as UAV crashes.
-To enhance the reliability of LLM in translating natural language tasks into operation code, techniques such as prompt engineering [13] and few-shot learning [14] have been widely adopted when designing LLM-driven operation code generation research for robots, including UAVs, and have demonstrated promising results.
-However, a set of existing approaches [2, 3, 4, 5] adopt an open-loop design that lacks mechanisms to verify the correctness of the operation code generated by LLMs.
-Such open-loop designs can leave questions such as cumulative error and temporal action consistency unsolved [15], and hence affect the reliability of their generated code, especially when handling complex tasks.
-Recently, closed-loop designs have been proposed in LLM-driven robotic operation code generation [6, 7, 8, 9, 10, 11], which enable evaluation of generated code and provide feedback for refinement, thereby steering robot execution towards its task objectives.
+To enhance the reliability of LLM in translating natural language tasks into operation code, techniques such as prompt engineering [PromptEngineering] and few-shot learning [few-shot] have been widely adopted when designing LLM-driven operation code generation research for robots, including UAVs, and have demonstrated promising results.
+However, a set of existing approaches [ChatGPTRobotics, TypeFly, CLEAR, GSCE] adopt an open-loop design that lacks mechanisms to verify the correctness of the operation code generated by LLMs.
+Such open-loop designs can leave questions such as cumulative error and temporal action consistency unsolved [OpenLoopQuestions], and hence affect the reliability of their generated code, especially when handling complex tasks.
+Recently, closed-loop designs have been proposed in LLM-driven robotic operation code generation [incoro, InteractivePlanning, CoPAL, CAPE, SelfRefine, trust], which enable evaluation of generated code and provide feedback for refinement, thereby steering robot execution towards its task objectives.
 Nevertheless, these approaches still face limitations from different aspects, including:
-  • Using numerical state observations for evaluation: approaches proposed in ref [6, 7] use numerical state observations in scenarios for evaluation and feedback generation.
-Due to the relatively limited numerical reasoning and processing capabilities of LLMs [16], an LLM-driven evaluation can overlook the latent semantics of robot actions and fail to achieve a comprehensive assessment of its movements and dynamics, leading to unreliable code refinement feedback.
-  • Physical execution for evaluation: closed-loop approaches, such as ref [8, 9, 7], employ physical execution of LLM-generated code to generate state observations for evaluation and code refinement.
-Although these approaches can be acceptable for robots that are not likely to cause physical damage during the refinement, their adoption in the context of UAV operation becomes risky because the physical execution of UAVs with code that potentially contains errors for refinement can easily lead to UAV crashes and safety concerns [17].
-  • Self-refinement using a single LLM: existing closed-loop approaches [10, 11] that a single LLM can be used for both code generation, evaluation, and refinement.
-However, self-evaluation may amplify its errors and lead to degradation in task success rate [18], which is also validated in the experimental evaluation of this work.
+  • Using numerical state observations for evaluation: approaches proposed in ref [incoro, InteractivePlanning] use numerical state observations in scenarios for evaluation and feedback generation.
+Due to the relatively limited numerical reasoning and processing capabilities of LLMs [NumberUnderstanding], an LLM-driven evaluation can overlook the latent semantics of robot actions and fail to achieve a comprehensive assessment of its movements and dynamics, leading to unreliable code refinement feedback.
+  • Physical execution for evaluation: closed-loop approaches, such as ref [CoPAL, CAPE, InteractivePlanning], employ physical execution of LLM-generated code to generate state observations for evaluation and code refinement.
+Although these approaches can be acceptable for robots that are not likely to cause physical damage during the refinement, their adoption in the context of UAV operation becomes risky because the physical execution of UAVs with code that potentially contains errors for refinement can easily lead to UAV crashes and safety concerns [uav-crash].
+  • Self-refinement using a single LLM: existing closed-loop approaches [SelfRefine, trust] that a single LLM can be used for both code generation, evaluation, and refinement.
+However, self-evaluation may amplify its errors and lead to degradation in task success rate [LLM-bias], which is also validated in the experimental evaluation of this work.
 To overcome these limitations, this paper proposes a novel closed-loop framework to further enhance the reliability of LLM-driven UAV operation code generation.
@@
 This feature enables efficient refinement for UAV operation code generation, and more importantly, eliminates the safety risks raised by potential incorrect code execution during refinement with physical UAV execution.
-Align with existing research [6, 7, 8, 9, 10, 11], this paper also focuses on using LLMs to generate operation code for deployment and task execution with the complete task description provided, but not real-time interactive UAV controls.
+Align with existing research [incoro, InteractivePlanning, CoPAL, CAPE, SelfRefine, trust], this paper also focuses on using LLMs to generate operation code for deployment and task execution with the complete task description provided, but not real-time interactive UAV controls.
 The high-level design of our framework is illustrated in Fig.
@@
 In our framework, the LLM configured as the Code Generator synthesizes or refines UAV operation code based on task description or evaluation feedback, and the LLM configured as the Evaluator assesses code performance and provides feedback to guide code refinement.
-The process unfolds as follows: 1) The Code Generator first produces an initial code based on the task description; 2) The code is then execute in simulation to produce raw numerical state observation, which are subsequently transformed into a semantic observations; 3) The Evaluator evaluates the semantic observations against the original task specification and returns evaluation feedback; 4) If the evaluation confirms that the observations has met the task objectives, the code will deploy on the UAV for task execution.
+The process unfolds as follows: 1) The Code Generator first produces an initial code based on the task description; 2) The code is then executed in simulation to produce raw numerical state observation, which are subsequently transformed into semantic observations; 3) The Evaluator evaluates the semantic observations against the original task specification and returns evaluation feedback; 4) If the evaluation confirms that the observations have met the task objectives, the code will deploy on the UAV for task execution.
 Otherwise, the evaluation feedback is returned to the Code Generator for iterative refinement.
@@
 For advanced tasks with a higher complexity, our framework achieves a $98.5\%$ completeness and an $85\%$ success rate, i.e., $98.5\%$ of required actions in all evaluated tasks are completed correctly and $85\%$ of tasks are entirely completed without any error.
-Compared with baseline approaches [5, 10, 7], our framework greatly improves both success rate and completeness when the task complexity increases.
+Compared with baseline approaches [GSCE, SelfRefine, InteractivePlanning], our framework greatly improves both success rate and completeness when the task complexity increases.
 Figure 2: Illustration of LLM-driven closed-loop UAV Operation and Refinement.
@@
 II-A LLM for Robotics
-Continual advancements in LLMs [19, 20] have demonstrated their exceptional capabilities in semantic understanding and context generation, making them promising when integrating with various robotics tasks.
-Research works from Microsoft [2] and Google [21] demonstrated how publicly accessible LLMs can be leveraged to generate operation code for various robotic tasks, including UAVs.
-Since then, a set of research works [3, 4, 5, 6, 7, 8, 9, 10, 11] have been proposed to enhance the quality and reliability of LLM-driven robotic tasks, especially with the integration of prompt engineering techniques [13] that help steer LLMs’ behavior and align their outputs with the tasks’ objectives.
-In ref [3], an end-to-end LLM framework is proposed that enables semantic planning with applications in quadcopter drones.
+Continual advancements in LLMs [deepseek, o3mini] have demonstrated their exceptional capabilities in semantic understanding and context generation, making them promising when integrating with various robotics tasks.
+Research works from Microsoft [ChatGPTRobotics] and Google [CodeasPolicies] demonstrated how publicly accessible LLMs can be leveraged to generate operation code for various robotic tasks, including UAVs.
+Since then, a set of research works [TypeFly, CLEAR, GSCE, incoro, InteractivePlanning, CoPAL, CAPE, SelfRefine, trust] have been proposed to enhance the quality and reliability of LLM-driven robotic tasks, especially with the integration of prompt engineering techniques [PromptEngineering] that help steer LLMs’ behavior and align their outputs with the tasks’ objectives.
+In ref [TypeFly], an end-to-end LLM framework is proposed that enables semantic planning with applications in quadcopter drones.
 This paper also introduced a specific programming language (MiniSpec) to improve the code generation efficiency of LLMs.
-In ref [4], a prompt-engineered robot control solution is proposed that synthesizes input from both LLMs and humans.
+In ref [CLEAR], a prompt-engineered robot control solution is proposed that synthesizes input from both LLMs and humans.
 This work aims to have the robotic system behavior programmed by emphasizing prompting LLMs and avoiding manipulating code.
-Recently, a prompt framework with enhanced reasoning, named GSCE [5] was proposed to better support LLM-driven UAV control.
+Recently, a prompt framework with enhanced reasoning, named GSCE [GSCE], was proposed to better support LLM-driven UAV control.
 GSCE integrates domain-specific guidelines, constraints, and examples into its system prompts when configuring the LLM, and hence improves its reliability of constraint-compliant code generation.
 While these methods all improve the LLM-driven code generation for robotic tasks from different aspects, they all adopt an open-loop design framework without error correction and adjustment for the generated code.
-Such a design can lead to potential cumulative errors and temporal action consistency issues [15] in the code generation.
+Such a design can lead to potential cumulative errors and temporal action consistency issues [OpenLoopQuestions] in the code generation.
 To address the concerns from the open-loop design, different closed-loop refinement strategies are increasingly adopted by recent research towards LLM-driven code generation for robotic tasks.
-In ref [6, 7], numerical observation of the robot’s execution status, such as actuator reading, sensor reading, and system information (e.g., location), is utilized to construct the feedback for LLMs.
+In ref [incoro, InteractivePlanning], numerical observation of the robot’s execution status, such as actuator reading, sensor reading, and system information (e.g., location), is utilized to construct the feedback for LLMs.
 These numerical observations-based feedback can positively contribute to the quality of LLM-generated operation code, however, their contribution is restricted when trying to assess the latent semantics of robot actions for a comprehensive assessment.
-This is because LLMs have relatively limited numerical reasoning and processing capabilities [16].
+This is because LLMs have relatively limited numerical reasoning and processing capabilities [NumberUnderstanding].
 In addition, to obtain accurate observations and leverage them for the refinement of LLM-generated code, executing the code with physical experiments is employed for feasible robotic tasks.
-For example, CoPAL [8] employs physical error feedback and run-time error feedback to refine the code when errors occur during physical execution.
-In ref [9], precondition errors are integrated into a prompt that aims to repair plans when the robot fails skill execution.
+For example, CoPAL [CoPAL] employs physical error feedback and run-time error feedback to refine the code when errors occur during physical execution.
+In ref [CAPE], precondition errors are integrated into a prompt that aims to repair plans when the robot fails skill execution.
 While evaluating code with physical experiments can provide accurate observations of errors and misalignment between LLM-generated code and task objectives that enable the construction of an effective refinement plan, it does not fit when moving to the context of many UAV tasks.
-Physical execution of code that may contain errors, especially before any refinement, has a high chance of causing the UAV to crash [17].
+Physical execution of code that may contain errors, especially before any refinement, has a high chance of causing the UAV to crash [uav-crash].
 This not only increases the hardware cost for the experiment but also can lead to safety concerns.
@@
 In recent research, other strategies to further improve the performance of closed-loop design have also been explored.
-For example, ref [22] utilizes human feedback to revise the code when misalignment in the generated plan is detected.
+For example, ref [preemptive] utilizes human feedback to revise the code when misalignment in the generated plan is detected.
 Although human feedback provides high-quality input for refinement, it limits the automation of the system and the performance depends on the expertise of the user.
-In addition, existing research [6, 7, 8, 9] separates the configuration of code generator and evaluator on different LLMs.
-This design aims to minimize the amplification of errors and bias [18] that can be caused by using a single LLM for both code generation and evaluation in previous research [10, 11].
+In addition, existing research [incoro, InteractivePlanning, CoPAL, CAPE] separates the configuration of the code generator and evaluator on different LLMs.
+This design aims to minimize the amplification of errors and bias [LLM-bias] that can be caused by using a single LLM for both code generation and evaluation in previous research [SelfRefine, trust].
 In our framework, we also use two separate LLMs and have customized configurations to improve their capabilities as the code generator and evaluator, respectively.
 II-B Prompt Engineering
-Recent studies have increasingly employed prompt engineering [13] to improve the reliability of LLM-driven robotic systems.
-For instance, GSCE proposes a prompt framework as a system prompt to improve the reliability of LLM-driven UAV control code generation [5].
-Prompt engineering typically leverages two core techniques: in-context learning [23] and Chain-of-Thought (CoT) reasoning [24].
+Recent studies have increasingly employed prompt engineering [PromptEngineering] to improve the reliability of LLM-driven robotic systems.
+For instance, GSCE proposes a prompt framework as a system prompt to improve the reliability of LLM-driven UAV control code generation [GSCE].
+Prompt engineering typically leverages two core techniques: in-context learning [URIAL] and Chain-of-Thought (CoT) reasoning [CoT].
 In-context learning enables the LLM to identify patterns from a limited set of examples and apply them to new situations.
-This facilitates the generation of control code that adheres to robot policy and how to ground task descriptions from a few examples [21, 25].
+This facilitates the generation of control code that adheres to robot policy and how to ground task descriptions from a few examples [CodeasPolicies, PromptBook].
 On the other hand, CoT prompts the LLM to articulate intermediate inference steps, which is valuable for robotic tasks requiring sequential, stepwise decision-making.
 By decomposing a task into steps, CoT encourages the production of code that aligns with each stage of the intended action plan.
-Prior studies have embedded CoT within the examples to guide LLMs through reasoning step-by-step via in-context learning [26, 25].
+Prior studies have embedded CoT within the examples to guide LLMs through reasoning step-by-step via in-context learning [ISR-LLM, PromptBook].
 In this study, we integrate both CoT and in-context learning strategies to enhance the quality of code generation and evaluation.
@@
 Given a UAV operation task description, the framework proposed in this paper aims to use LLMs to reason through the task description and generate accurate operation code after observation, evaluation, feedback, and refinement.
-Specifically, we formulate our problem by extending the LLM-driven UAV operation problem in [5] as $P=\langle s_{0},S,A,T,G\rangle$, where $s_{0}$ is the initial state, $S=(s_{0},s_{1},s_{2},\dots,s_{n})$ is the finite set of discrete states encoding the UAV’s spatial position and orientation, $A$ is the finite action (UAV skills) set, $T$ is the deterministic transition, and $G\subseteq S$ is the goal state.
+Specifically, we formulate our problem by extending the LLM-driven UAV operation problem in [GSCE] as $P=\langle s_{0},S,A,T,G\rangle$, where $s_{0}$ is the initial state, $S=(s_{0},s_{1},s_{2},\dots,s_{n})$ is the finite set of discrete states encoding the UAV’s spatial position and orientation, $A$ is the finite action (UAV skills) set, $T$ is the deterministic transition, and $G\subseteq S$ is the goal state.
 For a current state $S$, an action $a\in A$ is selected from the skill set $A$, leading to a deterministic state transition: $T:S\times A\rightarrow S$ that determines the next state.
@@
 The state transitions of $l$ yield state observations $O$, and evaluation $E=\langle O,P\rangle$ assesses whether $O$ satisfies the problem objectives and outputs $F$ as feedback.
-If $F$ is “YES”, the sequence of actions $l$ is deemed a valid solution solution that flies the UAV from $s_{0}$ to the goal state $(s_{n}\equiv G$); otherwise, $l$ is iteratively refined based on the feedback $F$.
+If $F$ is “YES”, the sequence of actions $l$ is deemed a valid solution that flies the UAV from $s_{0}$ to the goal state $(s_{n}\equiv G$); otherwise, $l$ is iteratively refined based on the feedback $F$.
 Figure 3: Overall design of LLM-driven closed-loop UAV operation code generation with semantic observations.
@@
 The LLM-driven code generation in our framework is achieved by configuring a general-purpose LLM agent (code generator) with UAV operation-related system prompts.
-We adopt prompt engineering strategies proposed by the GSCE framework [5], which have been demonstrated to be effective for enhancing the reasoning capabilities of LLMs in basic UAV operation code generation.
+We adopt prompt engineering strategies proposed by the GSCE framework [GSCE], which have been demonstrated to be effective for enhancing the reasoning capabilities of LLMs in basic UAV operation code generation.
 Specifically, the system prompts comprise guidelines, skill APIs, constraints, and examples.
@@
 It embeds the observations of the UAV’s historical trajectory dynamics, thereby enabling the evaluator to identify its deviations from the task objectives.
-Rather than getting trajectory observations from real robots [7, 8] that can cause a UAV crash if the LLM-generated code contains errors, our framework adopts simulation-based execution to generate trajectory observations.
+Rather than getting trajectory observations from real robots [InteractivePlanning, CoPAL] that can cause a UAV crash if the LLM-generated code contains errors, our framework adopts simulation-based execution to generate trajectory observations.
 When the LLM-generated code is executed in simulation, it produces the UAV’s state observations for each action.
 These observations are numerical representations of the UAV’s states, including the UAV’s position and orientation.
-However, numerical observations offer limited semantic insight for an LLM-based evaluator because LLM lacks prior knowledge about interpreting the semantic meaning of the numerical state vector [16].
+However, numerical observations offer limited semantic insight for an LLM-based evaluator because LLM lacks prior knowledge about interpreting the semantic meaning of the numerical state vector [NumberUnderstanding].
 To bridge this gap, we transform each numerical state observation into a step-wise semantic description and aggregate these into a semantic trajectory observation.
@@
 The evaluator aims to identify deviations (if any) between the semantic trajectory observation and the task description, and then construct feedback based on the deviations.
-Instead of using the same LLM used for code generation for self-feedback evaluation, our design employs a separate LLM to prevent the same LLM from amplifying its bias [18].
+Instead of using the same LLM used for code generation for self-feedback evaluation, our design employs a separate LLM to prevent the same LLM from amplifying its bias [LLM-bias].
 To improve the accuracy of LLM’s evaluation, we adopt the design principles of the code generator by configuring the evaluator as an LLM agent with a structured system prompt.
@@
 The feedback and refinement loop continues iteratively until the evaluator outputs “YES”, indicating that $E$ confirms that the refined action sequence $l$ successfully solves the task problem $P$ and the UAV operation code is ready for deployment.
-In addition, if the maximum number of iterations is reached without a successful evaluation, the evaluator suggests human intervention.
+In addition, if the maximum number of iterations is reached without a successful evaluation, human supervision should be involved to address potential errors in the code.
 V Experiment and Evaluation
 V-A Experimental Setup
-In our experiment, the implementation of our proposed framework and comparison methods [5, 10, 7] uses the OpenAI “o3-mini” (o3-mini-2025-01-31) model [20] as the foundational LLM.
+In our experiment, the implementation of our proposed framework and comparison methods [GSCE, SelfRefine, InteractivePlanning] uses the OpenAI “o3-mini” (o3-mini-2025-01-31) model [o3mini] as the foundational LLM.
 Based on the “o3-mini”, two separate LLMs are configured as the code generator and evaluator in our framework.
-The widely used UAV simulator, “AirSim” [27], is tailored to support the simulation component of our framework for code execution and state observations.
+The widely used UAV simulator, “AirSim” [AirSim], is tailored to support the simulation component of our framework for code execution and state observations.
 The UAV model selected for experiment and evaluation is the default quadcopter in AirSim.
@@
 Each Basic task involves one or two intuitive actions to evaluate LLM’s few-stage planning and basic spatial reasoning (e.g., “Turn 90° clockwise, then fly 4 meters left in the drone’s body frame”).
-We use the Basic task dataset established by existing research [5], which contains 44 tasks in total.
+We use the Basic task dataset established by existing research [GSCE], which contains 44 tasks in total.
 Differently, the Advanced tasks in our experiment require complex reasoning about the UAV’s position and orientation dynamics and typically require 6-19 actions to reach the goal state.
@@
 The “High” tasks require flying complex patterns with orientation requirements (e.g., fly a figure 8 while orienting in the flying direction).
-The “Very High” increases the complexity by adding requirements to the patterns (e.g., inspect an area twice with the second inspection flies the path reversely).
-The complete list of both Basic and Advanced tasks is available at our project website \@footnotemark All tasks were manually validated in our experiment to avoid potential simulator-induced errors that could affect the evaluation process.
+The “Very High” increases the complexity by adding requirements to the patterns (e.g., inspect an area twice with the second inspection flying the path reversely).
+The complete list of both Basic and Advanced tasks is available at our project website 1.
+All tasks were manually validated in our experiment to avoid potential simulator-induced errors that could affect the evaluation process.
 TABLE I: Examples of Tasks in Advanced Task Dataset
@@
 Next, ascend another 5 meters and fly the square pattern in reverse order to examine the same area.”
-Baseline Methods for Comparison: Three recently proposed methods [5, 10, 7] are selected as the baselines for comparison with our proposed approach, including
-  • GSCE [5]: An open-loop approach in which the GSCE code generator synthesizes control code based on the task prompt, without any refinement.
-  • LLM Self-feedback (Self-Refine) [10]: A single LLM performs both code generation and self-refinement.
-The LLM analyzes semantic trajectory observations to detect deviations from the task and refines code based on its self-assessment.
-  • Feedback with Numerical State Observations (Numerical) [7]: This method introduces a secondary LLM serving as evaluator designed according to section IV-D.
-The evaluator detects deviations from numerical state observations and provides corrective guidance for the code generator for refinement.
+Baseline Methods for Comparison: Three recently proposed methods [GSCE, SelfRefine, InteractivePlanning] are selected as the baselines for comparison with our proposed approach, including
+  • GSCE [GSCE]: An open-loop approach implemented by configuring a general LLM as UAV operation code generation agent (GSCE) with the system prompt that contains guidelines, skill APIs, constraints, and examples.
+The GSCE is first initialized with the system prompt; upon receiving a task prompt, it synthesizes control code based on the task instructions without any refinement.
+  • LLM Self-feedback (Self-Refine) [SelfRefine]: A single LLM responsible for both code generation and refinement, which extends the GSCE method by self-evaluation and self-refinement of the code.
+Specifically, after the initial code is generated, it is executed in simulation to produce semantic trajectory observations, then the GSCE is prompted to evaluate these observations against the task description and iteratively refine the code based on its self-assessment.
+  • Feedback with Numerical State Observations (Numerical) [InteractivePlanning]: Similar to our method, this method introduces a secondary LLM that evaluates the UAV states.
+In implementation, the code generator (GSCE) generates the initial code, which is then executed in simulation to produce UAV state observations in numerical representations of UAV state changes (format is the same as ground truth in the section V-B).
+Then, a separate LLM is configured as the evaluator with roles, rules, and references (numerical trajectory exemplars).
+The evaluator analyzes these numerical observations against the task description and provides corrective guidance to the code generator for subsequent refinement.
 V-B Evaluation Metrics
@@
 We assume the UAV’s state is fully observable at all times, thereby ensuring that each transition can be precisely monitored and compared against ground truth.
+Tolerance: These metrics evaluate whether the generated code can correctly operate the UAV to perform the actions required by the task description.
+As AirSim simulation injects white noise in sensor measurements to emulate real-world sensing behavior [AirSim], there are offsets in the UAV state estimates produced by AirSim.
+To filter such noises that could affect trajectory matching with ground truth, we set a tolerance threshold of 1 meter for each positional parameter $[x,y,z]$ and 4 degrees for the Yaw $\theta$ in our implementation.
 V-C Result and Analysis
@@
 The average SR and completeness across Basic and Advanced tasks are summarized in Table II.
-The results reflect the average value of three repetitions of experiments to mitigate the impact of potential randomness of LLM’s output [28].
+The results reflect the average value of three repetitions of experiments to mitigate the impact of potential randomness of LLM’s output [LLMrandomness].
 TABLE II: SR and Completeness of Proposed Methods
-Basic Task Advanced Task SR Completeness SR Completeness GSCE [5] $\% 100.0100.0\%$ $66.7\%$ $88.9\%$ Self-Refine [10] $\% 100.0100.0\%$ $50.0\%$ $74.3\%$ Numerical [7] $95.5\%$ $97.3\%$ $73.3\%$ $92.4\%$ Ours $\% 100.0100.0\%$ $\mathbf{85.0\%}$ $\mathbf{98.5\%}$
+Basic Task Advanced Task SR Completeness SR Completeness GSCE [GSCE] $\% 100.0100.0\%$ $66.7\%$ $88.9\%$ Self-Refine [SelfRefine] $\% 100.0100.0\%$ $50.0\%$ $74.3\%$ Numerical [InteractivePlanning] $95.5\%$ $97.3\%$ $73.3\%$ $92.4\%$ Ours $\% 100.0100.0\%$ $\mathbf{85.0\%}$ $\mathbf{98.5\%}$
 For the Basic tasks, all methods showed remarkable performance.
-It is notable that the Numerical method [7] has a slightly lower performance with a $95.5\%$ SR and a $97.3\%$ completeness.
+It is notable that the Numerical method [InteractivePlanning] has a slightly lower performance with a $95.5\%$ SR and a $97.3\%$ completeness.
 This indicates that relying on numerical state observations can lead to misinterpretation in evaluation, even when handling simple tasks.
@@
 While the impact of these factors is relatively negligible when handling basic tasks, their limitations are amplified when the task complexity increases, and affect the logical reasoning capabilities of LLMs to generate operation code correctly.
-Figure 6: Failure counts of different methods for each advanced task at different complexity levels
+TABLE III: Standard Deviation and $95\%$ Confidence Interval of Completeness for Tasks in Advanced Task Dataset
+Task GSCE [GSCE] Self-Refine [SelfRefine] Numerical [InteractivePlanning] Ours
+Standard Confidence Standard Confidence Standard Confidence Standard Confidence
+Deviation Interval Deviation Interval Deviation Interval Deviation Interval
+1 0 $[1.0,1.0]$ $0.385$ $[0.333,1.0]$ 0 $[1.0,1.0]$ 0 $[1.0,1.0]$
+2 $0.128$ $[0.778,1.0]$ $0.449$ $[0.222,1.0]$ $0.421$ $[0.222,1.0]$ 0 $[1.0,1.0]$
+3 0 $[1.0,1.0]$ $0.449$ $[0.222,1.0]$ 0 $[1.0,1.0]$ 0 $[1.0,1.0]$
+4 0 $[1.0,1.0]$ 0 $[1.0,1.0]$ 0 $[1.0,1.0]$ 0 $[1.0,1.0]$
+5 0 $[1.0,1.0]$ 0 $[1.0,1.0]$ 0 $[1.0,1.0]$ 0 $[1.0,1.0]$
+6 $0.173$ $[0.700,1.0]$ $0.058$ $[0.9,1.0]$ $0.115$ $[0.8,1.0]$ 0 $[1.0,1.0]$
+7 $0.167$ $[0.444,0.778]$ $0.321$ $[0.444,1.0]$ $0.280$ $[0.444,1.0]$ 0 $[1.0,1.0]$
+8 $0.096$ $[0.778,0.944]$ $0.064$ $[0.889,1.0]$ $0.032$ $[0.944,1.0]$ $0.192$ $[0.611,0.944]$
+9 $0.238$ $[0.588,1.0]$ $0.068$ $[0.412,0.529]$ 0 $[1.0,1.0]$ 0 $[1.0,1.0]$
+10 $0.513$ $[0.111,1.0]$ 0 $[0.944,0.944]$ $0.289$ $[0.5,1.0]$ 0 $[1.0,1.0]$
+11 0 $[1.0,1.0]$ 0 $[0.2,0.2]$ 0 $[1.0,1.0]$ 0 $[1.0,1.0]$
+12 $0.052$ $[0.909,1.0]$ $0.472$ $[0.182,1.0]$ 0 $[1.0,1.0]$ 0 $[1.0,1.0]$
+13 0 $[1.0,1.0]$ $0.527$ $[0.059,1.0]$ 0 $[1.0,1.0]$ 0 $[1.0,1.0]$
+14 0 $[1.0,1.0]$ $0.064$ $[0.889,1.0]$ 0 $[1.0,1.0]$ 0 $[1.0,1.0]$
+15 0 $[1.0,1.0]$ $0.289$ $[0.5,1.0]$ 0 $[1.0,1.0]$ 0 $[1.0,1.0]$
+16 $0.102$ $[0.8,1.0]$ $0.346$ $[0.4,1.0]$ 0 $[1.0,1.0]$ 0 $[1.0,1.0]$
+17 $0.058$ $[0.9,1.0]$ 0 $[0.4,0.4]$ $0.173$ $[0.6,0.9]$ $0.058$ $[0.9,1.0]$
+18 $0.229$ $[0.579,1.0]$ 0 $[1.0,1.0]$ $0.169$ $[0.684,1.0]$ $0.030$ $[0.947,1.0]$
+19 0 $[1.0,1.0]$ $0.272$ $[0.529,1.0]$ $0.238$ $[0.588,1.0]$ 0 $[1.0,1.0]$
+20 0 $[0.944,0.944]$ 0 $[0.722,0.722]$ $0.032$ $[0.944,1.0]$ 0 $[0.944,0.944]$
+Overall $0.164$ $[0.843,0.926]$ $0.325$ $[0.66,0.822]$ $0.178$ $[0.876,0.964]$ $\mathbf{0.055}$ $\mathbf{[0.969,0.995]}$
+Statistical Analysis of Completeness: To evaluate the stability and consistency of our method, we performed statistical analysis for the completeness of advanced tasks.
+Specifically, the standard deviation and the $95\%$ confidence interval are measured for each advanced task across all three repetitions.
+As summarized in Table III, our method has a low overall standard deviation of $0.055$ and a high overall confidence interval $[0.969,0.995]$.
+These values also remain stable among almost all advanced tasks except for one of the failure cases.
+A more detailed analysis of the failure cases is also provided below.
+Compared with the baseline methods, which have higher standard deviations and wider confidence intervals for most tasks as shown in Table III, our method clearly achieves better stability and consistency.
+Figure 6: Failure counts of different methods for each advanced task at different complexity levels.
 Figure 7: Completeness of each failed task of different methods at different complexity levels
-TABLE III: Failure Cases Completeness in Each Level of Complexity
+TABLE IV: Failure Cases Completeness in Each Level of Complexity
 Low Medium High Very Overall
 High
-GSCE [5] $90.9\%$ $81.9\%$ $0.7\% 77$$6\% 77.$$7.4\%$
-Self-Refine [10] $21.6\%$ $41.1\%$ $67.4\%$ $57.8\%$ $48.5\%$
-Numerical [7] $62.6\%$ $67.8\%$ $77.2\%$ $69.5\%$
+GSCE [GSCE] $90.9\%$ $81.9\%$ $0.7\% 77$$6\% 77.$$7.4\%$
+Self-Refine [SelfRefine] $21.6\%$ $41.1\%$ $67.4\%$ $57.8\%$ $48.5\%$
+Numerical [InteractivePlanning] $62.6\%$ $67.8\%$ $77.2\%$ $69.5\%$
 Ours $\mathbf{90.0\%}$ $\mathbf{94.4\%}$ $\mathbf{90.4\%}$ $\mathbf{91.6\%}$
@@
 In particular, we aim to understand how many actions are correctly completed when a task fails.
-As summarized in Table III, our method achieves a high average completeness of $91.6\%$ for all failed tasks, i.e., the operation code generated by our method can still complete $91.6\%$ of the required actions in the task and remains highly aligned with the task objectives even when errors occur.
+As summarized in Table IV, our method achieves a high average completeness of $91.6\%$ for all failed tasks, i.e., the operation code generated by our method can still complete $91.6\%$ of the required actions in the task and remains highly aligned with the task objectives even when errors occur.
 Differently, the average completeness of the compared baseline methods ranges from $48.5\%$ to $77.4\%$ for their failed tasks, which indicates errors occur for a significant portion of the required actions in these tasks.
@@
 Our method has failures on only four tasks during three rounds of experiments.
-As a comparison, the baseline methods failed 11 (GSCE [5]), 17 (Self-Refine [10]), and 9 (Numerical [7]) tasks, respectively.
-In addition, Table III and Fig.
+As a comparison, the baseline methods failed 11 (GSCE [GSCE]), 17 (Self-Refine [SelfRefine]), and 9 (Numerical [InteractivePlanning]) tasks, respectively.
+In addition, Table IV and Fig.
 7 also show that our method has a stable high completeness of $90\%+$ for all failed tasks at different complexity levels.
@@
 Based on our evaluation results, our method achieves a higher reliability compared with the baseline methods, not only because of its high confidence in generating fully correct operation code, but also its high completeness of the required actions of tasks when a small amount of errors occur.
-TABLE IV: Evaluator Configurations Description and Example System Prompts
+TABLE V: Evaluator Configurations Description and Example System Prompts
 Evaluator Configuration Description
@@
 In our proposed framework, the operation code generated by the Code Generator LLM will be refined through multiple iterations.
-The increased rounds of iterations provide additional opportunities to improve the code quality, however, they also introduce additional system cost (e.g., LLM API cost and processing time).
+The increased rounds of iterations provide additional opportunities to improve the code quality.
+However, they also introduce additional system cost (e.g., LLM API cost and processing time).
 To balance this tradeoff and identify an appropriate number of iterations, we performed experiments of our framework on advanced tasks with 0 to $10$ rounds of iterations.
-Our results presented in Fig.
-8 show the changes of SR and completeness with the different number of iterations.
-Specifically, both SR and completeness are boosted after the iteration is introduced for refinement, and then have a slowly increasing trend until the $6\textsuperscript{th}$ iteration.
-After that, the performance of the subsequent iterations becomes relatively stable in terms of both SR and completeness, and even declines slightly.
-This is because LLMs can exhibit sycophantic behavior and tend to adapt or change outputs to appear agreeable [29, 30], and thus may perform unnecessary changes after sufficient refinement has been performed.
-Our detailed examination identified that such a situation occurred on task 7 in our experiment.
+As described in Algorithm 1, the early stopping is used in our design when the evaluator outputs “YES”, which prevents additional unnecessary refinement iterations.
+The mean of our three repetition results is presented in Fig.
+8, which shows the changes of SR and completeness with the increment of refinement iterations.
+Specifically, the mean iteration for our method to start converging with a relatively stable performance is 3.
+With a mean iteration of 6, our method is converged with the best performance in terms of both completeness and success rate.
+The performance declines slightly after the $6\textsuperscript{th}$ iteration because LLMs can exhibit sycophantic behavior and tend to adapt or change outputs to appear agreeable [flipflop, LLM-behaviors] when receiving a negative query, and thus may perform unnecessary changes after sufficient refinement has been performed.
 This observation demonstrates that iterative refinement can effectively contribute to the LLM-driven UAV operation code generation.
-Meanwhile, it is also necessary to impose an upper bound on refinement iterations to balance the tradeoff.
-According to our evaluation results, we set the maximum number of iterations to $6$.
-However, this number may need to be adjusted for different types of tasks.
+Meanwhile, it is also necessary to impose an upper bound on refinement iterations to balance the tradeoff between the success rate and system cost.
+As presented in Fig.
+9, the computational cost and OpenAI API usage cost increase with the number of iterations.
+Thus, it is unnecessary to continue the refinement when the gain of success rate becomes very limited, i.e., after $6$ iterations in our setting.
+The number of iterations can be adjusted for different types of tasks.
 For example, setting a lower number when the tasks to complete have a relatively low complexity.
 Figure 8: Curves of SR (left) and completeness (right) over the number of refinement iterations for our method on advanced tasks.
+Figure 9: Curves of per-task time and price over the number of refinement iterations.
 V-D Evaluator Configuration Analysis
@@
 For each configuration, we instantiate the evaluator’s system prompt by including only the indicated components.
-The detailed description of each evaluator configuration with sample system prompts for roles, rules, and references is provided in Table IV.
-TABLE V: Performance over Evaluator Configurations
+The detailed description of each evaluator configuration with sample system prompts for roles, rules, and references is provided in Table V.
+TABLE VI: Performance over Evaluator Configurations
 (1) (2) (3) (4)